### v1

# On the numerical stability of sketched GMRES

Abstract
We perform a backward stability analysis of preconditioned sketched GMRES [Nakatsukasa and Tropp, SIAM J. Matrix Anal. Appl, 2024] for solving linear systems $Ax=b$, and show that the backward stability at iteration $i$ depends on the conditioning of the Krylov basis $B_{1:i}$ as long as the condition number of $AB_{1:i}$ can be bounded by $1/O(\bm{u})$, where $\bm{u}$ is the unit roundoff. Under this condition, we show that sketched GMRES is backward stable as long as the condition number of $B_{1:i}$ is not too large. Under additional assumptions, we then show that the stability of a restarted implementation of sketched GMRES can be independent of the condition number of $B_{1:i}$, and restarted sketched GMRES is backward stable. We also derive sharper bounds that explain why the backward error can be small even in cases when the basis $B_{1:i}$ is very ill-conditioned, which has been observed in the literature but not yet explained theoretically. We present numerical experiments to demonstrate the conclusions of our analysis, and also show that adaptively restarting where appropriate allows us to recover backward stability in sketched GMRES.
Keywords: GMRES, sketched GMRES, randmonized algorithms
AMS subject classifications (2020). 65F10, 65F50, 65G50
1 Introduction
In this paper, we analyze the numerical stability of sketched GMRES (sGMRES) [10] for solving nonsymmetric and nonsingular linear systems of the form
$Ax=b,\enspace A\in\mathbb{R}^{n\times n}$ (1)
with a given matrix $A$ and a right-hand side $b$.
Although standard GMRES [12] is one of the most popular methods to solve (1), it comes at the expense of constructing a matrix $V_{1:i}\in\mathbb{R}^{n\times i}$ with columns that form an orthonormal basis for the Krylov subspace
$\mathcal{K}_{i}(A,r^{(0)})=\operatorname{span}\{r^{(0)},Ar^{(0)},\dotsc,A^{i-1}r^{(0)}\},$
where $i$ is the iteration number, and $r^{(0)}=b-Ax^{(0)}$ is the initial residual corresponding to an initial approximation $x^{(0)}$.
Sketched GMRES was proposed in [10] as an attractive alternative to GMRES that exploits the technique of randomized sketching [9] to avoid working with an orthonormal basis. Instead, the basis is only partially orthogonalized, thereby reducing the arithmetic cost from $O(ni^{2})$ to $O(i^{3}+ni\log i)$.
It has been observed in numerical experiments in [10] that sGMRES can fail once the condition number of the basis $B_{1:i}$ becomes too large, and methods such as the sketch-and-select Arnoldi algorithm [7] have been proposed as a means to ensure the condition number of the basis $B_{1:i}$ does not grow too rapidly. Indeed, numerical experiments in [7] showed that working with a well-conditioned basis improved the overall convergence behavior of sketched GMRES.
In this paper, we rigorously analyze the backward stability of preconditioned sketched GMRES, and justify the experimental observations made in the literature [10, 7]. Namely, we show that the backward error of sketched GMRES depends on $\kappa(Z_{1:i})$ as long as $\kappa(AZ_{1:i})$ is not too large, where $Z_{1:i}=M_{R}^{-1}B_{1:i}$ with the right preconditioner $M_{R}$. We then provide a backward stability result for a restarted version of sketched GMRES, which under certain assumptions, allows us to recover the numerical stability of sketched GMRES when $\kappa(Z_{1:i})$ grows large, suggesting we should monitor the conditioning of the basis throughout the iterations and restart when necessary. We present numerical experiments to demonstrate the numerical behavior we expect to observe in sketched GMRES as predicted by our new analysis, and demonstrate the improvements in the backward error that can be obtained from the adaptive restart approach.
[…]
$\min\{\epsilon:(A+\Delta A)\hat{x}=b+\Delta b,\bigl{\|}\Delta A\bigr{\rVert}_{\mathsf{F}}\leq\epsilon\bigl{\|}A\bigr{\rVert}_{\mathsf{F}},\|\Delta b\rVert\leq\epsilon\|b\rVert\}=\frac{\|b-A\hat{x}\rVert}{\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\|\hat{x}\rVert+\|b\rVert}.$
The remainder of this paper is organized as follows. In Section 2, we give an overview of GMRES and the MOD-GMRES framework [3] for performing a backward stability analysis of different GMRES variants. We then discuss preconditioned sketched GMRES in Section 3, and analyze the backward stability of preconditioned sketched GMRES and restarted sketched GMRES, respectively, in Sections 4 and 5. Numerical experiments are presented in Section 6.
2 Preconditioned GMRES
In this section, we briefly outline the basics of preconditioned GMRES with a left preconditioner $M_{L}$ for solving linear systems of the form (1). Beginning with an initial solution approximation $x^{(0)}$ and associated initial residual $r^{(0)}=b-Ax^{(0)}$, preconditioned GMRES iteratively generates a matrix $V_{1:i}\in\mathbb{R}^{n\times i}$ with columns that form an orthonormal basis for the Krylov subspace $\mathcal{K}_{i}(M_{L}^{-1}A,M_{L}^{-1}r^{(0)})$. The orthonormal basis $V_{1:i}$ is typically constructed via the Arnoldi process, and satisfies the Arnoldi relation
$M_{L}^{-1}AV_{1:i}=V_{1:(i+1)}\underline{H_{i}},$ (2)
where $\underline{H_{i}}\in\mathbb{R}^{(i+1)\times i}$ is an upper Hessenberg matrix that stores the coefficients of the orthogonalization. The orthonormality of $V_{1:i}$ then allows for the GMRES least squares solution,
$t^{(i)}=\operatorname*{\arg\min}_{t\in\mathcal{K}_{i}(M_{L}^{-1}A,M_{L}^{-1}r^{(0)})}\|M_{L}^{-1}r^{(0)}-M_{L}^{-1}At\rVert,$ (3)
to be computed by solving a small $(i+1)\times i$ least squares problem,
$y^{(i)}=\operatorname*{\arg\min}_{y\in\mathbb{R}^{i}}\|M_{L}^{-1}r^{(0)}-M_{L}^{-1}AV_{1:i}y^{(i)}\rVert=\operatorname*{\arg\min}_{y\in\mathbb{R}^{i}}\|\|M_{L}^{-1}r^{(0)}\|e_{1}-\underline{H_{1:i}}y\rVert,$
and setting $t^{(i)}=V_{1:i}y^{(i)}$. The solution approximation is then updated via $x^{(i)}=x^{(0)}+t^{(i)}$. If a right preconditioner $M_{R}$ is also taken into account, it can be integrated with the basis $Z_{1:i}$ such that $Z_{1:i}=M_{R}^{-1}V_{1:i}$.
There are many algorithmic variants of GMRES arising from different implementation options, such as the choice of preconditioner (if any), or the choice of orthogonalization method used within the Arnoldi procedure. Furthermore, for optimal performance it is common to exploit techniques such as deflation, mixed precision, or randomization within GMRES. In order to understand the finite precision behavior of any given GMRES variant, a backward error analysis framework was developed in [3], which simplifies the process of deriving bounds for the attainable normwise backward and forward errors of the computed solutions of different GMRES variants. The framework is built from a backward error analysis of a generic GMRES algorithm known as modular GMRES or MOD-GMRES, that is composed of three elemental operations:
  1. 1.
    Generate the preconditioned Krylov basis $Z_{1:i}=M_{R}^{-1}V_{1:i}$ with the Krylov basis $V_{1:i}$, and compute $W_{1:i}=M_{L}^{-1}AZ_{1:i}$.
  2. 2.
    Solve the least squares problem $y^{(i)}=\operatorname*{\arg\min}_{y}\|\tilde{b}-W_{1:i}y\rVert$ with $\tilde{b}=M_{L}^{-1}b$.
  3. 3.
    Compute the approximate solution $x^{(i)}=Z_{1:i}y^{(i)}$.
Without loss of generality, we assume $x^{(0)}=0$ in this framework and our analysis of sGMRES in Section 4. By specializing these operations and meeting their assumptions, as outlined in the framework, modular GMRES can describe many GMRES implementations and variants, and can be used to perform their backward error analysis. For example, well-known results from [5] were derived in [3] using the framework. In particular, it was shown that GMRES with Householder orthogonalization (HH-GMRES) produces a computed solution whose backward error is on the order of the unit roundoff, and is thus backward stable. Similarly, the backward stability result for GMRES with modified Gram-Schmidt (MGS-GMRES), derived in [11], was also recovered using the framework in [3].
3 Preconditioned Sketched GMRES
Sketched GMRES aims to avoid the $O(ni^{2})$ arithmetic cost of performing an expensive full orthogonalization of the basis vectors in standard GMRES. Using the technique of randomized sketching, the GMRES minimization problem (3) is replaced by the sketched problem
$y^{(i)}=\operatorname*{\arg\min}_{y\in\mathbb{R}^{i}}\|SM_{L}^{-1}r^{(0)}-SM_{L}^{-1}AZ_{1:i}y\rVert,$ (4)
for a preconditioned basis $Z_{1:i}=M_{R}^{-1}B_{1:i}$ with a (non-orthogonal) Krylov basis $B_{1:i}$ and sketching operator $S\in\mathbb{R}^{s\times n}$ with $i<s\ll n$. The sketched problem (4) can then be constructed and solved inexpensively, even without orthonormality of the basis $B_{1:i}$. The solution of the sketched problem serves as a good approximation of the original problem as long as the matrix $S$ is chosen to be an $\epsilon$-subspace embedding for the Krylov subspace $\mathcal{K}_{i}(M_{L}^{-1}AM_{R}^{-1},M_{L}^{-1}r^{(0)})$, i.e., $S$ satisfies
$(1-\epsilon)\|v\|^{2}\leq\|Sv\|^{2}\leq(1+\epsilon)\|v\|^{2},$ (5)
for $\epsilon\in(0,1)$, and every $v\in\mathcal{K}_{i}(M_{L}^{-1}AM_{R}^{-1},M_{L}^{-1}r^{(0)})$. The sketching matrix $S$ needs to be drawn at random to ensure the relation (5) holds with a high probability. In practice, the sketched least squares problem (4) is computed by solving a triangular linear system
$T_{1:i}y^{(i)}=U_{1:i}^{\top}(SM_{L}^{-1}r^{(0)}),$ (6)
where the orthonormal matrix $U_{1:i}$ and the triangular matrix $T_{1:i}$ come from the QR factorization of the reduced matrix $C_{1:i}=SM_{L}^{-1}AZ_{1:i}$, i.e., $C_{1:i}=U_{1:i}T_{1:i}$.
A detailed outline of one cycle of preconditioned sketched GMRES is presented in Algorithm 1, and a restarted version is given in Algorithm 2.
0:  A matrix $A\in\mathbb{R}^{n\times n}$, a right-hand side $b\in\mathbb{R}^{n}$, an initial approximation $x^{(0)}\in\mathbb{R}^{n}$, the maximal number of iterations $m$, a subspace embedding $S\in\mathbb{R}^{s\times n}$ ($s>m$), left and right preconditioners $M_{L}$ and $M_{R}$.
0:  A computed solution $x\in\mathbb{R}^{n}$ approximating the solution of $Ax=b$.
1:  $r^{(0)}\leftarrow b-Ax^{(0)}$ and $g\leftarrow SM_{L}^{-1}r^{(0)}$.
2:  $B_{1}\leftarrow r^{(0)}/\|r^{(0)}\rVert$.
3:  for $i=1:m$ do
4:     Generate the $i$-th column $B_{i}$ of the basis by Arnoldi process, and $Z_{i}\leftarrow M_{R}^{-1}B_{i}$.
5:     $W_{i}\leftarrow M_{L}^{-1}AZ_{i}$.
6:     Sketch reduced matrix $C_{1:i}\leftarrow SW_{1:i}$.
7:     Compute the QR factorization $C_{1:i}=U_{1:i}T_{1:i}$ using $C_{1:i-1}=U_{1:i-1}T_{1:i-1}$, where $U_{1:i}$ is an orthonormal matrix and $T_{1:i}$ is an upper triangular matrix.
8:     if the stopping criterion is satisfied then
9:        Solve the triangular system $T_{1:i}y^{(i)}=U_{1:i}^{\top}g$ to obtain $y^{(i)}$.
10:        return  $x=x^{(i)}\leftarrow x^{(0)}+Z_{1:i}y^{(i)}$.
11:     end if
12:  end for
Algorithm 1 Preconditioned sketched GMRES (sGMRES)
0:  A matrix $A\in\mathbb{R}^{n\times n}$, a right-hand side $b\in\mathbb{R}^{n}$, an initial approximation $x^{(0)}\in\mathbb{R}^{n}$, the maximal number of iterations $m$, a subspace embedding $S\in\mathbb{R}^{s\times n}$ ($s>m$), left and right preconditioners $M_{L}$ and $M_{R}$, and the maximum number of restarts nrestarts.
0:  A computed solution $x\in\mathbb{R}^{n}$ approximating the solution of $Ax=b$.
1:  for j = 1:nrestarts do
2:     Call preconditioned sGMRES with input $M_{L}$, $M_{R}$, $A$, $b$, and $x^{(0)}$ to obtain $x$.
3:     Update the initial approximate solution for the next iteration, $x^{(0)}\leftarrow x$.
4:  end for
Algorithm 2 Restarted preconditioned sketched GMRES
It is clear from experimental observations made in [10, 7] that the basis condition number should not grow too large throughout the iteration in order to ensure numerical stability of sketched GMRES. The non-orthogonal basis can be constructed using the $t$ truncated Arnoldi process, where each new Krylov vector is orthogonalized against the previous $t$ vectors. Alternatively, the sketch-and-select method [7] uses the sketched basis $SB_{i}$ to select $t$ vectors, among all previously computed vectors, with which to perform the partial orthogonalization, with the aim of generating a basis with smaller condition number than that generated by truncated Arnoldi.
4 Backward stability of preconditioned sketched GMRES
In this section, we aim to analyze the backward stability of preconditioned sketched GMRES (Algorithm 1) by primarily following the analysis of the MOD-GMRES framework [3]. The first $i$ iterations of preconditioned sketched GMRES can be summarized in the following steps:
  1. 1.
    Generate the preconditioned Krylov basis $Z_{1:i}=M_{R}^{-1}B_{1:i}$ with the (non-orthogonal) Krylov basis $B_{1:i}$, and compute $W_{1:i}=M_{L}^{-1}AZ_{1:i}$.
  2. 2.
    Solve the least squares problem $y^{(i)}=\operatorname*{\arg\min}_{y}\|Sb-SW_{1:i}y\rVert$ by solving a triangular linear system (6) using the QR factorization of $C_{1:i}=SW_{1:i}$.
  3. 3.
    Compute the approximate solution $x^{(i)}=Z_{1:i}y^{(i)}$.
Notice that the Krylov basis $B_{1:i}$ is non-orthogonal and may even be ill-conditioned, and the least squares problems solved in Step 2 are different for sketched GMRES and MOD-GMRES.
The main result of our backward error analysis relies on various assumptions on each of these steps, which we list here.
Step 1: Matrix–matrix product
For the computed preconditioned basis $\hat{Z}_{1:i}\in\mathbb{R}^{n\times i}$, we assume that the computed matrix–matrix product $\hat{W}_{1:i}$ computed by Line 5 in Algorithm 1 satisfies
$\hat{W}_{1:i}=M_{L}^{-1}A\hat{Z}_{1:i}+\Delta W_{1:i},\quad\bigl{\|}\Delta W_{1:i}\bigr{\rVert}_{\mathsf{F}}\leq\varepsilon_{W}\|M_{L}^{-1}\rVert\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\bigl{\|}\hat{Z}_{1:i}\bigr{\rVert}_{\mathsf{F}},$ (7)
with $\varepsilon_{W}\leq O(\bm{u})$ from the standard rounding error analysis of [8]. Similarly, we assume that the computed result of $\tilde{b}:=M_{L}^{-1}b$ computed by Line 1 in Algorithm 1 satisfies
$\hat{b}=\tilde{b}+\Delta b,\quad\|\Delta b\rVert\leq\varepsilon_{b}\|M_{L}^{-1}\rVert\|b\rVert.$ (8)
Note that we assume $x^{(0)}=0$ in the analysis, and thus $r^{(0)}=b$.
Step 2: Least squares solver
In iteratively solving the least squares problem in Step 2, we assume that there is a special dimension $k\leq n$, known as the key dimension [3], for which the computed solution is guaranteed to have reached a small backward error, i.e., $\hat{b}$ lies in the range of non-singular basis $\hat{W}_{1:k}$. The definition of the key dimension is as follows:
$\sigma_{\min}(\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix})$ $\leq\omega_{\kappa(r,W)}\bigl{\|}\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix}\bigr{\rVert}_{\mathsf{F}},$ (9)
$\sigma_{\min}(\hat{W}_{1:k})$ $\geq\omega_{\kappa(W)}\bigl{\|}\hat{W}_{1:k}\bigr{\rVert}_{\mathsf{F}}$ (10)
for all $\phi>0$, where $\omega_{\kappa(r,W)}$ and $\omega_{\kappa(W)}$ are scalar functions of $n$, $k$, and $\bm{u}$, depending on the method for solving the least squares problem. If $\omega_{\kappa(r,W)}=O(\bm{u})$ and $\omega_{\kappa(W)}=O(\bm{u})$, (10) guarantees that $\hat{W}_{1:k}$ is a numerically non-singular basis of the subspace with dimension $k$ and then (9) means that $\hat{b}$ lies in the range of $\hat{W}_{1:k}$ numerically. Furthermore, it also implies that $\min_{y}\|\hat{b}-\hat{W}_{1:k}y\rVert$ is sufficiently small.
Using the notion of the key dimension, we now aim to write down an expression for the computed sketched residual vector at the $k^{th}$ iteration of sketched GMRES. We then provide a bound on this sketched residual in Lemma 1, and use this bound in the proof of our final backward stability result in Theorem 2.
We first set up the least squares problem by constructing the matrix $\hat{C}_{1:k}$, and the vector $g:=SM_{L}^{-1}b$. We assume that
$\hat{C}_{1:k}$ $=S\hat{W}_{1:k}+\Delta C_{1:k},\quad\|\Delta C_{j}\rVert\leq\varepsilon_{S}\|S\hat{W}_{j}\rVert,\quad\forall j\leq k,$ (11)
$\hat{g}$ $=S\hat{b}+\Delta g,\quad\|\Delta g\rVert\leq\varepsilon_{S}\|M_{L}^{-1}\rVert\|S\hat{b}\rVert,$ (12)
and furthermore, that the computed solution $\hat{y}^{(k)}$ of the least squares problem in Step 2 satisfies
$\begin{split}&\hat{y}^{(k)}=\operatorname*{\arg\min}_{y}\|\hat{g}+\Delta\hat{g}-(\hat{C}_{1:k}+\Delta\hat{C}_{1:k})y\rVert,\\ &\|\begin{bmatrix}\Delta\hat{g}&\Delta\hat{C}_{1:k}\end{bmatrix}e_{j}\rVert\leq\varepsilon_{sls}\|\begin{bmatrix}\hat{g}&\hat{C}_{1:k}\end{bmatrix}e_{j}\rVert,\quad\forall j\leq k+1,\end{split}$ (13)
with $\varepsilon_{sls}\leq O(\bm{u})$ from [8, Theorem 20.3]. Substituting $\hat{C}_{1:k}$ and $\hat{g}$ in (13) with (11) and (12), we have
$\begin{split}&\hat{y}^{(k)}=\operatorname*{\arg\min}_{y}\big{\|}S\tilde{\hat{b}}+\Delta\tilde{g}-(S\hat{W}_{1:k}+\Delta\tilde{C}_{1:k})y\big{\rVert},\\ &\|\Delta\tilde{g}\rVert\leq\|\Delta g\rVert+\|\Delta\hat{g}\rVert\leq(1+\varepsilon_{sls})\varepsilon_{S}\|S\hat{b}\rVert+\varepsilon_{sls}\|S\hat{b}\rVert\leq\bigl{(}\varepsilon_{sls}+(1+\varepsilon_{sls})\varepsilon_{S}\bigr{)}\|S\hat{b}\rVert,\\ &\|\Delta\tilde{C}_{j}\rVert\leq\|\Delta C_{j}\rVert+\|\Delta\hat{C}_{j}\rVert\leq\bigl{(}\varepsilon_{sls}+(1+\varepsilon_{sls})\varepsilon_{S}\bigr{)}\|S\hat{W}_{j}\rVert,\quad\forall j\leq k.\end{split}$ (14)
We can now prove a bound on the norm of the sketched residual at iteration $k$, i.e., the norm of the vector
$r^{(k)}_{S}:=S\hat{b}+\Delta\tilde{g}-(S\hat{W}_{1:k}+\Delta\tilde{C}_{1:k})\hat{y}^{(k)}.$ (15)
Lemma 1.
Assume that $\hat{W}_{1:k}$ satisfies
$\sigma_{\min}(\begin{bmatrix}\phi S\hat{b}&S\hat{W}_{1:k}\end{bmatrix})$ $\leq c_{1}(n,k)\bigl{(}\varepsilon_{S}+\varepsilon_{sls}\bigr{)}\bigl{\|}\begin{bmatrix}\phi S\hat{b}&S\hat{W}_{1:k}\end{bmatrix}\bigr{\rVert}_{\mathsf{F}},$ (16)
$\sigma_{\min}(S\hat{W}_{1:k})$ $\geq c_{2}(n,k)\bigl{(}\varepsilon_{S}+\varepsilon_{sls}\bigr{)}\bigl{\|}S\hat{W}_{1:k}\bigr{\rVert}_{\mathsf{F}},$ (17)
where $S$ is an $\epsilon$-subspace embedding for the space spanned by the matrix $\begin{bmatrix}\hat{b}&\hat{W}_{1:k}\end{bmatrix}$ satisfying (5), and $c_{1}(n,k)$ and $c_{2}(n,k)$ are low degree polynomials related to $n$ and $k$. Also assume that $\hat{y}^{(k)}$ satisfies (14). Then
$\|r^{(k)}_{S}\rVert\leq c(n,k)\bigl{(}\varepsilon_{S}+\varepsilon_{sls}\bigr{)}\bigl{\|}S\hat{W}_{1:k}\bigr{\rVert}_{\mathsf{F}}\|\hat{y}^{(k)}\rVert,$ (18)
where $c(n,k)$ is also a low degree polynomial related to $n$ and $k$.
Proof.
We follow the proof in [3] and employ [11, Theorem 2.4] to bound $\|r^{(k)}_{S}\rVert$. Applying [11, Theorem 2.4] to the least squares problem (14), we have
$\|r^{(k)}_{S}\rVert^{2}\leq\sigma_{\min}^{2}\bigl{(}\begin{bmatrix}\phi(S\hat{b}+\Delta\tilde{g})&S\hat{W}_{1:k}+\Delta\tilde{C}_{1:k}\end{bmatrix}\bigr{)}\biggl{(}\frac{1}{\phi^{2}}+\frac{\|\hat{y}^{(k)}\rVert^{2}}{1-\delta_{k}^{2}}\biggr{)},$ (19)
with $\delta_{k}$ defined by
$\delta_{k}=\frac{\sigma_{\min}\bigl{(}\begin{bmatrix}\phi(S\hat{b}+\Delta\tilde{g})&S\hat{W}_{1:k}+\Delta\tilde{C}_{1:k}\end{bmatrix}\bigr{)}}{\sigma_{\min}\bigl{(}S\hat{W}_{1:k}+\Delta\tilde{C}_{1:k}\bigr{)}}<1.$ (20)
We choose $\phi$ such that $\phi^{-2}=\|\hat{y}^{(k)}\rVert^{2}/(1-\delta_{k}^{2})$ as in [3, (3.17)], which can simplify (19) as
$\|r^{(k)}_{S}\rVert\leq\sqrt{2}\,\sigma_{\min}\bigl{(}\begin{bmatrix}\phi(S\hat{b}+\Delta\tilde{g})&S\hat{W}_{1:k}+\Delta\tilde{C}_{1:k}\end{bmatrix}\bigr{)}\phi^{-1}.$ (21)
Then we will show that there exists a $\phi>0$ satisfying $\phi^{-2}=\|\hat{y}^{(k)}\rVert^{2}/(1-\delta_{k}^{2})$. This is equivalent to showing that there exists a $\phi>0$ such that $f(\phi)=0$, where the function $f(\phi)$ is defined as
$\begin{split}f(\phi)&=\sigma_{\min}^{2}\bigl{(}S\hat{W}_{1:k}+\Delta\tilde{C}_{1:k}\bigr{)}-\sigma_{\min}^{2}\bigl{(}\begin{bmatrix}\phi(S\hat{b}+\Delta\tilde{g})&S\hat{W}_{1:k}+\Delta\tilde{C}_{1:k}\end{bmatrix}\bigr{)}\\ &\quad-\|\hat{y}^{(k)}\rVert^{2}\phi^{2}\sigma_{\min}^{2}\bigl{(}S\hat{W}_{1:k}+\Delta\tilde{C}_{1:k}\bigr{)}.\end{split}$ (22)
Notice that $f(0)>0$ and $f(\|\hat{y}^{(k)}\rVert^{-1})<0$. By continuity, this implies that there exists $\phi\in(0,\|\hat{y}^{(k)}\rVert^{-1})$ such that $\phi^{-2}=\|\hat{y}^{(k)}\rVert^{2}/(1-\delta_{k}^{2})$ and $\delta_{k}<1$.
It remains to bound $\sigma_{\min}\bigl{(}\begin{bmatrix}\phi(S\hat{b}+\Delta\tilde{g})&S\hat{W}_{1:k}+\Delta\tilde{C}_{1:k}\end{bmatrix}\bigr{)}$ and $\phi$ in (21). By the assumption (16), and the estimation of $\|\Delta\tilde{g}\rVert$ and $\|\Delta\tilde{C}_{1:k}\rVert$ shown in (14), we have
$\begin{split}\sigma_{\min}\bigl{(}&\begin{bmatrix}\phi(S\hat{b}+\Delta\tilde{g})&S\hat{W}_{1:k}+\Delta\tilde{C}_{1:k}\end{bmatrix}\bigr{)}\\ &\leq c_{1}(n,k)\bigl{(}\varepsilon_{S}+\varepsilon_{sls}\bigr{)}\bigl{\|}\begin{bmatrix}\phi S\hat{b}&S\hat{W}_{1:k}\end{bmatrix}\bigr{\rVert}_{\mathsf{F}}\\ &\leq c_{1}(n,k)\bigl{(}\varepsilon_{S}+\varepsilon_{sls}\bigr{)}\bigl{(}\|S\hat{b}\rVert\phi+\bigl{\|}S\hat{W}_{1:k}\bigr{\rVert}_{\mathsf{F}}\bigr{)}.\end{split}$ (23)
To establish a bound for $\phi$, it suffices to bound $\delta_{k}$. Combining (23) with the assumption (17) and the estimation (14) of $\|\Delta\tilde{C}_{1:k}\rVert$, $\delta_{k}$ can be bounded by
$\begin{split}\delta_{k}\leq\frac{c_{1}(n,k)\bigl{(}\varepsilon_{S}+\varepsilon_{sls}\bigr{)}\bigl{(}\|S\hat{b}\rVert\phi+\bigl{\|}S\hat{W}_{1:k}\bigr{\rVert}_{\mathsf{F}}\bigr{)}}{c_{2}(n,k)\bigl{(}\varepsilon_{S}+\varepsilon_{sls}\bigr{)}\bigl{\|}S\hat{W}_{1:k}\bigr{\rVert}_{\mathsf{F}}}.\end{split}$ (24)
We now show that $\|S\hat{b}\rVert\phi$ can be bounded by $\bigl{\|}S\hat{W}_{1:k}\bigr{\rVert}_{\mathsf{F}}$. From the definition of $r^{(k)}_{S}$ in (15), we have
$\|S\hat{b}\rVert\phi\leq\|r^{(k)}_{S}\rVert\phi+\|\Delta\tilde{g}\rVert\phi+\bigl{\|}S\hat{W}_{1:k}\bigr{\rVert}_{\mathsf{F}}\|\hat{y}^{(k)}\rVert\phi+\|\Delta\tilde{C}_{1:k}\rVert\|\hat{y}^{(k)}\rVert\phi.$
Notice that $\|\hat{y}^{(k)}\rVert\phi\leq 1$ since $\phi\in(0,\|\hat{y}^{(k)}\rVert^{-1})$. Furthermore, using (21), (23), and the estimation of $\|\Delta\tilde{g}\rVert$ and $\|\Delta\tilde{C}_{1:k}\rVert$ from (14), we obtain
$\|S\hat{b}\rVert\phi\leq\tilde{c}_{1}(n,k)(\varepsilon_{sls}+\varepsilon_{S})\bigl{(}\|S\hat{b}\rVert\phi+\bigl{\|}S\hat{W}_{1:k}\bigr{\rVert}_{\mathsf{F}}\bigr{)}+\bigl{\|}S\hat{W}_{1:k}\bigr{\rVert}_{\mathsf{F}},$
which means
$\|S\hat{b}\rVert\phi\leq\frac{1+\tilde{c}_{1}(n,k)(\varepsilon_{sls}+\varepsilon_{S})}{1-\tilde{c}_{1}(n,k)(\varepsilon_{sls}+\varepsilon_{S})}\bigl{\|}S\hat{W}_{1:k}\bigr{\rVert}_{\mathsf{F}},$ (25)
and then $\delta_{k}\leq 1/2$ by choosing appropriate $c_{1}(n,k)$ and $c_{2}(n,k)$ in (24). Thus, we have
$\phi=\frac{\|\hat{y}^{(k)}\rVert}{\sqrt{1-\delta_{k}}}\leq 2\|\hat{y}^{(k)}\rVert,$ (26)
and, using (25), (26), (23), and (21), we can conclude the proof. ∎
Lemma 1 provides the values of $\omega_{\kappa(r,W)}$ and $\omega_{\kappa(W)}$ for defining the key dimension, that is, $c_{1}(n,k)\bigl{(}\varepsilon_{S}+\varepsilon_{sls}\bigr{)}$ and $c_{2}(n,k)\bigl{(}\varepsilon_{S}+\varepsilon_{sls}\bigr{)}$, respectively.
Step 3: Computation of the approximate solution
For $\hat{Z}_{1:k}\in\mathbb{R}^{n\times k}$ and $\hat{y}^{(k)}\in\mathbb{R}^{k}$, we assume that
$\hat{x}^{(k)}=\hat{Z}_{1:k}\hat{y}^{(k)}+\Delta x^{(k)},\quad\|\Delta x^{(k)}\rVert\leq\varepsilon_{x}\|\hat{Z}_{1:k}\rVert\|\hat{y}^{(k)}\rVert$ (27)
with $\varepsilon_{x}\leq O(\bm{u})$ from [8].
We are now ready to describe the backward stability of preconditioned sketched GMRES (Algorithm 1) in the following theorem.
Theorem 1.
Assume that Algorithm 1 is applied with a computed preconditioned basis $\hat{Z}_{1:k}\in\mathbb{R}^{n\times k}$ and an $\epsilon$-subspace embedding $S\in\mathbb{R}^{s\times n}$, where (7), (8), (11), (12), (13), and (27) are satisfied. Then there exists $k\leq n$ such that $c_{1}(n,k)\bigl{(}\varepsilon_{S}+\varepsilon_{sls}\bigr{)}\sqrt{\frac{1-\epsilon}{1+\epsilon}}\kappa(\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix})\geq 1$. If it further holds that
$\sqrt{k}\,c_{2}(n,k)\bigl{(}\varepsilon_{S}+\varepsilon_{sls}\bigr{)}\sqrt{\frac{1+\epsilon}{1-\epsilon}}\kappa(\hat{W}_{1:k})\leq 1/2,\qquad\varepsilon_{x}\kappa(\hat{Z}_{1:k})\leq 1/2,$ (28)
then, with high probability,
$\begin{split}\lVert b-&A\hat{x}^{(k)}\rVert\\ &\leq\tilde{c}(n,k)\biggl{(}\frac{1+\epsilon}{1-\epsilon}\bigl{(}\varepsilon_{S}+\varepsilon_{sls}\bigr{)}+\varepsilon_{W}+\varepsilon_{x}\biggr{)}\kappa(M_{L})\kappa(\hat{Z}_{1:k})\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\|\hat{x}^{(k)}\rVert\\ &\quad+\biggl{(}\frac{1+\epsilon}{1-\epsilon}(1+\varepsilon_{b})\bigl{(}\varepsilon_{sls}+(1+\varepsilon_{sls})\varepsilon_{S}\bigr{)}+\varepsilon_{b}\biggr{)}\kappa(M_{L})\|b\rVert\end{split}$ (29)
with a low degree polynomial $\tilde{c}(n,k)$ related to $n$ and $k$. Furthermore, if $\varepsilon_{W}$, $\varepsilon_{b}$, $\varepsilon_{sls}$, $\varepsilon_{S}$, $\varepsilon_{x}\leq O(\bm{u})$, then
$\|b-A\hat{x}^{(k)}\rVert\leq\frac{1+\epsilon}{1-\epsilon}O(\bm{u})\kappa(M_{L})\bigl{(}\kappa(\hat{Z}_{1:k})\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\|\hat{x}^{(k)}\rVert+\|b\rVert\bigr{)}.$ (30)
Proof.
The residual of $\hat{x}^{(k)}$ can be bounded by the preconditioned residual as follows:
$\|b-A\hat{x}^{(k)}\rVert=\|M_{L}\bigl{(}M_{L}^{-1}b-M_{L}^{-1}A\hat{x}^{(k)})\bigr{\rVert}\leq\|M_{L}\rVert\|M_{L}^{-1}b-M_{L}^{-1}A\hat{x}^{(k)}\rVert,$ (31)
which means that we only need to focus on the preconditioned residual $\|M_{L}^{-1}b-M_{L}^{-1}A\hat{x}^{(k)}\rVert$. By substituting $\hat{x}^{(k)}$ with (27), $M_{L}^{-1}b$ with (8), and $M_{L}^{-1}A\hat{Z}_{1:k}$ with (7), we have
$\|M_{L}^{-1}b-M_{L}^{-1}A\hat{x}^{(k)}\rVert\leq\|\hat{b}-\hat{W}_{1:k}\hat{y}^{(k)}\rVert+\|\Delta W_{1:k}\hat{y}^{(k)}\rVert+\|\Delta b\rVert+\|M_{L}^{-1}\rVert\|A\Delta x^{(k)}\rVert.$ (32)
Furthermore, from [10, Fact 2.2] and the definition (15) of $r^{(k)}_{S}$, we have
$\begin{split}\|\hat{b}-\hat{W}_{1:k}\hat{y}^{(k)}\rVert&\leq\frac{1}{1-\epsilon}\|S\hat{b}-S\hat{W}_{1:k}\hat{y}^{(k)}\rVert\\ &\leq\frac{1}{1-\epsilon}\|r^{(k)}_{S}\rVert+\frac{1}{1-\epsilon}\bigl{(}\|\Delta\tilde{g}\rVert+\bigl{\|}\Delta\tilde{C}_{1:k}\bigr{\rVert}_{\mathsf{F}}\|\hat{y}^{(k)}\rVert\bigr{)}.\end{split}$ (33)
Combining (33) with (32), it follows that
$\begin{split}\|M_{L}^{-1}\hat{b}-M_{L}^{-1}A\hat{x}^{(k)}\rVert&\leq\frac{1}{1-\epsilon}\|r^{(k)}_{S}\rVert+\frac{1}{1-\epsilon}\bigl{(}\|\Delta\tilde{g}\rVert+\bigl{\|}\Delta\tilde{C}_{1:k}\bigr{\rVert}_{\mathsf{F}}\|\hat{y}^{(k)}\rVert\bigr{)}\\ &\quad+\|\Delta W_{1:k}\rVert\|\hat{y}^{(k)}\rVert+\|\Delta b\rVert+\|M_{L}^{-1}\rVert\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\|\Delta x^{(k)}\rVert.\end{split}$ (34)
It remains to bound $\|r^{(k)}_{S}\rVert$ and $\|\hat{y}^{(k)}\rVert$. To apply Lemma 1 to bound $\|r^{(k)}_{S}\rVert$, we first need to verify (16) and (17). From [1, Corollary 2.2], the condition numbers of the sketched matrices satisfy
$\kappa(S\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix})\geq\sqrt{\frac{1-\epsilon}{1+\epsilon}}\kappa(\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix})\quad\text{and}\quad\kappa(S\hat{W}_{1:k})\leq\sqrt{\frac{1+\epsilon}{1-\epsilon}}\kappa(\hat{W}_{1:k}),$
which implies, from $c_{1}(n,k)\bigl{(}\varepsilon_{S}+\varepsilon_{sls}\bigr{)}\sqrt{\frac{1-\epsilon}{1+\epsilon}}\kappa(\begin{bmatrix}\phi b&\hat{W}_{1:k}\end{bmatrix})\geq 1$ and (28),
$\sigma_{\min}(S\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix})$ $\leq c_{1}(n,k)\bigl{(}\varepsilon_{S}+\varepsilon_{sls}\bigr{)}\|S\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix}\rVert$
$\leq c_{1}(n,k)\bigl{(}\varepsilon_{S}+\varepsilon_{sls}\bigr{)}\bigl{\|}S\begin{bmatrix}\phi b&\hat{W}_{1:k}\end{bmatrix}\bigr{\rVert}_{\mathsf{F}},$
$\sqrt{k}\,c_{2}(n,k)\bigl{(}\varepsilon_{S}+\varepsilon_{sls}\bigr{)}\kappa(S\hat{W}_{1:k})$ $\leq\sqrt{k}\,c_{2}(n,k)\bigl{(}\varepsilon_{S}+\varepsilon_{sls}\bigr{)}\sqrt{\frac{1+\epsilon}{1-\epsilon}}\kappa(\hat{W}_{1:k})\leq 1,$
which guarantees that (16) and (17) hold. Thus, from Lemma 1, we obtain the bound on $\|r^{(k)}_{S}\rVert$:
$\|r^{(k)}_{S}\rVert\leq c(n,k)\bigl{(}\varepsilon_{S}+\varepsilon_{sls}\bigr{)}(1+\epsilon)\bigl{\|}\hat{W}_{1:k}\bigr{\rVert}_{\mathsf{F}}\|\hat{y}^{(k)}\rVert.$ (35)
Using the bounds on $\|r^{(k)}_{S}\rVert$, i.e., (35), as well as the bounds on $\Delta\tilde{g}$, $\Delta\tilde{C}_{1:k}$, $\Delta W_{1:k}$, $\Delta b$, and $\Delta x^{(k)}$ shown in (14), (7), (8), and (27), then (34) can be bounded as
$\begin{split}\|M_{L}^{-1}\hat{b}-M_{L}^{-1}A\hat{x}^{(k)}\rVert&\leq\frac{1+\epsilon}{1-\epsilon}c(n,k)\bigl{(}\varepsilon_{S}+\varepsilon_{sls}\bigr{)}\bigl{\|}\hat{W}_{1:k}\bigr{\rVert}_{\mathsf{F}}\|\hat{y}^{(k)}\rVert\\ &\quad+\frac{1+\epsilon}{1-\epsilon}\bigl{(}\varepsilon_{sls}+(1+\varepsilon_{sls})\varepsilon_{S}\bigr{)}\biggl{(}\|\hat{b}\rVert+\bigl{\|}\hat{W}_{1:k}\bigr{\rVert}_{\mathsf{F}}\|\hat{y}^{(k)}\rVert\biggr{)}\\ &\quad+\varepsilon_{W}\|M_{L}^{-1}\rVert\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\bigl{\|}\hat{Z}_{1:k}\bigr{\rVert}_{\mathsf{F}}\|\hat{y}^{(k)}\rVert+\varepsilon_{b}\|M_{L}^{-1}\rVert\|b\rVert\\ &\quad+\varepsilon_{x}\|M_{L}^{-1}\rVert\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\bigl{\|}\hat{Z}_{1:k}\bigr{\rVert}_{\mathsf{F}}\|\hat{y}^{(k)}\rVert.\end{split}$ (36)
Furthermore, noticing that $\bigl{\|}\hat{W}_{1:k}\bigr{\rVert}_{\mathsf{F}}\leq(1+\varepsilon_{W})\|M_{L}^{-1}\rVert\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\bigl{\|}\hat{Z}_{1:k}\bigr{\rVert}_{\mathsf{F}}$ from (7), (36) can be simplified as
$\begin{split}\lVert&M_{L}^{-1}\hat{b}-M_{L}^{-1}A\hat{x}^{(k)}\rVert\\ &\leq\biggl{(}\frac{1+\epsilon}{1-\epsilon}c(n,k)(1+\varepsilon_{W})\bigl{(}\varepsilon_{S}+\varepsilon_{sls}\bigr{)}+\sqrt{k}\,\varepsilon_{W}+\sqrt{k}\,\varepsilon_{x}\biggr{)}\|M_{L}^{-1}\rVert\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\|\hat{Z}_{1:k}\rVert\|\hat{y}^{(k)}\rVert\\ &\quad+\biggl{(}\frac{1+\epsilon}{1-\epsilon}(1+\varepsilon_{b})\bigl{(}\varepsilon_{sls}+(1+\varepsilon_{sls})\varepsilon_{S}\bigr{)}+\varepsilon_{b}\biggr{)}\|M_{L}^{-1}\rVert\|b\rVert.\end{split}$ (37)
Thus, using (27) we have
$\|\hat{y}^{(k)}\rVert\leq\frac{\|\hat{x}^{(k)}\rVert}{\sigma_{\min}(\hat{Z}_{1:k})-\varepsilon_{x}\|\hat{Z}_{1:k}\rVert}.$ (38)
Combining (37) with (38), we derive
$\begin{split}\lVert&M_{L}^{-1}\hat{b}-M_{L}^{-1}A\hat{x}^{(k)}\rVert\\ &\leq\biggl{(}\frac{1+\epsilon}{1-\epsilon}c(n,k)(1+\varepsilon_{W})\bigl{(}\varepsilon_{S}+\varepsilon_{sls}\bigr{)}+\sqrt{k}\,\varepsilon_{W}+\sqrt{k}\,\varepsilon_{x}\biggr{)}\frac{\|M_{L}^{-1}\rVert\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\|\hat{Z}_{1:k}\rVert\|\hat{x}^{(k)}\rVert}{\sigma_{\min}(\hat{Z}_{1:k})-\varepsilon_{x}\|\hat{Z}_{1:k}\rVert}\\ &\quad+\biggl{(}\frac{1+\epsilon}{1-\epsilon}(1+\varepsilon_{b})\bigl{(}\varepsilon_{sls}+(1+\varepsilon_{sls})\varepsilon_{S}\bigr{)}+\varepsilon_{b}\biggr{)}\|M_{L}^{-1}\rVert\|b\rVert\\ &\leq\biggl{(}\frac{1+\epsilon}{1-\epsilon}c(n,k)(1+\varepsilon_{W})\bigl{(}\varepsilon_{S}+\varepsilon_{sls}\bigr{)}+\sqrt{k}\,\varepsilon_{W}+\sqrt{k}\,\varepsilon_{x}\biggr{)}\frac{\|M_{L}^{-1}\rVert\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\kappa(\hat{Z}_{1:k})\|\hat{x}^{(k)}\rVert}{1-\varepsilon_{x}\kappa(\hat{Z}_{1:k})}\\ &\quad+\biggl{(}\frac{1+\epsilon}{1-\epsilon}(1+\varepsilon_{b})\bigl{(}\varepsilon_{sls}+(1+\varepsilon_{sls})\varepsilon_{S}\bigr{)}+\varepsilon_{b}\biggr{)}\|M_{L}^{-1}\rVert\|b\rVert,\end{split}$
which, together with (31) and the assumption (28) concludes the proof. ∎
Remark 1.
If the preconditioners satisfy $M_{L}=M_{R}=I$, then the relative backward error can be bounded as
$\frac{\|b-A\hat{x}^{(k)}\rVert}{\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\|\hat{x}^{(k)}\rVert+\|b\rVert}\leq\frac{1+\epsilon}{1-\epsilon}O(\bm{u})\kappa(\hat{B}_{1:k}).$ (39)
Remark 2.
Let $\varepsilon_{*}$ denote all the $\varepsilon$ with any subscript mentioned in this section. Considering $\varepsilon_{*}\leq O(\bm{u})$, note that the assumption (28) in Theorem 1 can be replaced by
$O(\bm{u})\sqrt{\frac{1+\epsilon}{1-\epsilon}}\kappa(M_{L})\kappa(A)\kappa(\hat{B}_{1:k})\leq 1/2$
since
$\begin{split}\kappa(\hat{W}_{1:k})&\leq\frac{\|M_{L}^{-1}A\hat{Z}_{1:k}\rVert+O(\bm{u})\|M_{L}^{-1}\rVert\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\|\hat{B}_{1:k}\rVert}{\sigma_{\min}(M_{L}^{-1}A\hat{Z}_{1:k})-O(\bm{u})\|M_{L}^{-1}\rVert\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\|\hat{Z}_{1:k}\rVert}\\ &\leq 2\kappa(M_{L}^{-1}A\hat{Z}_{1:k})+O(\bm{u})\kappa(M_{L})\kappa(A)\kappa(\hat{Z}_{1:k})\\ &\leq(2+O(\bm{u}))\kappa(M_{L})\kappa(A)\kappa(\hat{Z}_{1:k})\end{split}$ (40)
from (7). Therefore, if $S$ is chosen to be a $n\times n$ identity matrix and $\hat{B}_{1:k}$ is chosen to be an orthonormal Krylov basis, Theorem 1 recovers a similar backward stability result as [3, Theorem 3.1].
Notice that $\kappa(\hat{W}_{1:k})\leq 3\kappa(M_{L}^{-1}AZ_{1:j})$ from (40) if $O(\bm{u})\kappa(M_{L})\kappa(A)\kappa(\hat{Z}_{1:k})\leq 1$. As a consequence of Theorem 1, we can conclude that if $\kappa(M_{L}^{-1}AZ_{1:j})$ is sufficiently small, then the backward stability of sGMRES depends on the condition number of the preconditioned basis $Z_{1:j}$. In particular, sGMRES is backward stable provided $\kappa(Z_{1:j})$ is not too large.
In has been observed in numerical experiments in [6, Figure 9] that in some cases, sGMRES will continue to converge even when $\kappa(\hat{Z}_{1:k})>10^{15}$, and so the bound shown in Theorem 1 is not tight. Therefore, we present the following lemma to show a tighter bound, which can be directly derived from (37) by noticing
$\|b-A\hat{x}^{(k)}\rVert\leq\|M_{L}\rVert\lVert M_{L}^{-1}\hat{b}-M_{L}^{-1}A\hat{x}^{(k)}\rVert\quad\text{and}\quad\|M_{L}\rVert\|M_{L}^{-1}\rVert=\kappa(M_{L}).$
Lemma 2.
Assume that Algorithm 1 is applied with a computed preconditioned basis $\hat{Z}_{1:k}\in\mathbb{R}^{n\times k}$ and an $\epsilon$-subspace embedding $S\in\mathbb{R}^{s\times n}$, where (7), (8), (11), (12), (13), and (27) are satisfied. Then there exists $k\leq n$ such that $c_{1}(n,k)\bigl{(}\varepsilon_{S}+\varepsilon_{sls}\bigr{)}\sqrt{\frac{1-\epsilon}{1+\epsilon}}\kappa(\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix})\geq 1$. If it further holds that
$\sqrt{k}\,c_{2}(n,k)\bigl{(}\varepsilon_{S}+\varepsilon_{sls}\bigr{)}\sqrt{\frac{1+\epsilon}{1-\epsilon}}\kappa(\hat{W}_{1:k})\leq 1/2,$
then, with high probability,
$\begin{split}\|b-A\hat{x}^{(k)}\rVert&\leq\tilde{c}(n,k)\biggl{(}\frac{1+\epsilon}{1-\epsilon}\bigl{(}\varepsilon_{S}+\varepsilon_{sls}\bigr{)}+\varepsilon_{W}+\varepsilon_{x}\biggr{)}\kappa(M_{L})\|\hat{Z}_{1:k}\rVert\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\|\hat{y}^{(k)}\rVert\\ &\quad+\biggl{(}\frac{1+\epsilon}{1-\epsilon}(1+\varepsilon_{b})\bigl{(}\varepsilon_{sls}+(1+\varepsilon_{sls})\varepsilon_{S}\bigr{)}+\varepsilon_{b}\biggr{)}\kappa(M_{L})\|b\rVert\end{split}$ (41)
with a low degree polynomial $\tilde{c}(n,k)$ related to $n$ and $k$. Furthermore, if $\varepsilon_{W}$, $\varepsilon_{b}$, $\varepsilon_{sls}$, $\varepsilon_{S}$, $\varepsilon_{x}\leq O(\bm{u})$, then
$\|b-A\hat{x}^{(k)}\rVert\leq\frac{1+\epsilon}{1-\epsilon}O(\bm{u})\kappa(M_{L})\bigl{(}\|\hat{Z}_{1:k}\rVert\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\|\hat{y}^{(k)}\rVert+\|b\rVert\bigr{)}.$ (42)
Remark 3.
Comparing the bounds given in (29) and (30) within Theorem 1, Lemma 2 utilizes $\|\hat{y}^{(k)}\rVert$ in the bounds (41) and (42) as an alternative to employing $\|\hat{x}^{(k)}\rVert/\sigma_{\min}(\hat{Z}_{1:k})$. Based on (38) and the presumption (28), it follows that
$\|\hat{y}^{(k)}\rVert\leq\frac{2\|\hat{x}^{(k)}\rVert}{\sigma_{\min}(\hat{Z}_{1:k})}.$
This implies that Lemma 2 offers a sharper bound compared to Theorem 1, as demonstrated by
$\kappa(M_{L})\bigl{(}\|\hat{Z}_{1:k}\rVert\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\|\hat{y}^{(k)}\rVert+\|b\rVert\bigr{)}\leq 2\kappa(M_{L})\bigl{(}\kappa(\hat{Z}_{1:k})\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\|\hat{x}^{(k)}\rVert+\|b\rVert\bigr{)}.$
Furthermore, in practice, $\|\hat{y}^{(k)}\rVert$ can be much smaller than $\|\hat{x}^{(k)}\rVert/\sigma_{\min}(\hat{Z}_{1:k})$. This implies that the relative backward error bound based on $\|\hat{y}^{(k)}\rVert$ can remain small or potentially decrease even when $\kappa(\hat{Z}_{1:k})>10^{15}$.
Remark 4.
By Lemma 2, the relative backward error can be bounded as
$\frac{\|b-A\hat{x}^{(k)}\rVert}{\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\|\hat{x}^{(k)}\rVert+\|b\rVert}\leq\frac{1+\epsilon}{1-\epsilon}O(\bm{u})\kappa(M_{L})\frac{\|\hat{Z}_{1:k}\rVert\|\hat{y}^{(k)}\rVert}{\|\hat{x}^{(k)}\rVert},$
and can be simplified under the condition $M_{L}=M_{R}=I$ to
$\frac{\|b-A\hat{x}^{(k)}\rVert}{\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\|\hat{x}^{(k)}\rVert+\|b\rVert}\leq\frac{1+\epsilon}{1-\epsilon}O(\bm{u})\frac{\|\hat{B}_{1:k}\rVert\|\hat{y}^{(k)}\rVert}{\|\hat{x}^{(k)}\rVert}.$ (43)
5 Adaptively restarting sGMRES
Similar to the case of standard GMRES, restarting can be regarded as an iterative refinement process, and can remove the influence of both $\kappa(M_{L})$ and $\kappa(Z_{1:j})$ on the backward stability of sGMRES, as shown in the following version of [3, Theorem 4.1] describing the stability of restarted sketched GMRES (Algorithm 2).
Theorem 2.
Assume that Algorithm 1 is restarted, such that for restart number $i\geq 1$, we compute a basis $\hat{Z}_{1:k^{(i)}}^{(i)}\in\mathbb{R}^{n\times k^{(i)}}$ and a randomized subspace embedding $S^{(i)}\in\mathbb{R}^{s\times n}$, where (7), (8), (11), (12), (13), and (27) are satisfied for $\varepsilon_{W}^{(i)}$, $\varepsilon_{sls}^{(i)}$, $\varepsilon_{S}^{(i)}$, $\varepsilon_{x}^{(i)}\leq O(\bm{u})$. Then there exists $k^{(i)}\leq n$ such that
$O(\bm{u})\sqrt{\frac{1-\epsilon}{1+\epsilon}}\kappa\bigl{(}\begin{bmatrix}\phi r^{(i)}&\hat{W}_{1:k^{(i)}}^{(i)}\end{bmatrix}\bigr{)}\geq 1.$
Assume that
$\Lambda^{(i)}_{1}=O(\bm{u})\frac{1+\epsilon}{1-\epsilon}\kappa(\hat{Z}_{1:k^{(i)}}^{(i)})\kappa_{\mathsf{F}}(A)\ll 1$
holds for $i$. Then, with high probability, the backward error is reduced in the $i$-th restart cycle by a factor $\Lambda^{(i)}_{1}$ until it holds that
$\frac{\|b-A\hat{x}\rVert}{\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\|\hat{x}\rVert+\|b\rVert}\leq O(\bm{u}).$ (44)
As a consequence of Theorem 2, it is clear that a restarted variant of sketched GMRES (Algorithm 2) may allow us to recover the stability of sGMRES when the conditioning of the basis grows large.
It is clear from our analysis that care should be taken to ensure $\kappa(Z_{1:i})$ does not get too large as the iteration progresses. In fact, as noted in [10], due to the subspace embedding property (5), the condition number of $Z_{1:i}$ can be monitored inexpensively by explicitly computing the condition number of $SZ_{1:i}$. The numerical stability of sGMRES can thus be controlled by monitoring the condition number of $SZ_{1:i}$ in every iteration, and restarting when it grows above some specified tolerance $\texttt{tol}_{B}$. For the case of a truncated Arnoldi implementation, or the sketch-and-select Arnoldi implementations, the truncation parameter $t$ can then be increased upon restart to ensure the condition number of the basis $Z_{1:i}$ does not grow as rapidly as it did in the previous cycle.
As an alternative, due to the result of Lemma 2, we may also opt to monitor
$\tau_{i}:=\frac{\|Z_{1:i}\rVert\|y^{(i)}\rVert}{\|x^{(i)}\rVert},$
from (43) at each iteration, since it provides a sharper bound on the backward error, and can be inexpensively estimated as
$\tilde{\tau}_{i}:=\frac{\|SZ_{1:i}\rVert\|y^{(i)}\rVert}{\|Sx^{(i)}\rVert}.$
We may then opt to restart when $\texttt{tol}_{\tau}\cdot\tau_{i}\geq 1$, for some chosen tolerance $\texttt{tol}_{\tau}$. Again, optionally, we may then wish to increase the value of the truncation parameter $t$. In Algorithm 3, we present a hybrid approach, where we monitor $\tilde{\tau}_{i}$, and if it is large, we then restart. Before restarting, we also check if $SB_{1:i}$ is large, and if it is, we increase the truncation parameter, by, e.g., doubling it.
0:  A matrix $A\in\mathbb{R}^{n\times n}$, a right-hand side $b\in\mathbb{R}^{n}$, an initial approximation $x^{(0)}\in\mathbb{R}^{n}$, the maximal number of iterations $m$, a subspace embedding $S\in\mathbb{R}^{s\times n}$ ($s>m$), left and right preconditioners $M_{L}$ and $M_{R}$, and truncation parameter $t$, $\texttt{tol}_{\tau}$ and $\texttt{tol}_{B}$ for the restarting strategy.
0:  A computed solution $x\in\mathbb{R}^{n}$ approximating the solution of $Ax=b$.
1:  $r^{(0)}\leftarrow b-Ax^{(0)}$ and $g\leftarrow SM_{L}^{-1}r^{(0)}$.
2:  $\tilde{x}^{(0)}\leftarrow Sx^{(0)}$.
3:  $B_{1}\leftarrow r^{(0)}/\|r^{(0)}\rVert$.
4:  for $i=1:m$ do
5:     Generate the $i$-th column $B_{i}$ of the basis by the Arnoldi process, and $Z_{i}\leftarrow M_{R}^{-1}B_{i}$.
6:     $W_{i}\leftarrow M_{L}^{-1}AZ_{i}$.
7:     Sketch reduced matrix $C_{1:i}\leftarrow SW_{1:i}$.
8:     Compute the QR factorization $C_{1:i}=U_{1:i}T_{1:i}$ using $C_{1:i-1}=U_{1:i-1}T_{1:i-1}$, where $U_{1:i}$ is an orthonormal matrix and $T_{1:i}$ is an upper triangular matrix.
9:     Solve the triangular system $T_{1:i}y^{(i)}=U_{1:i}^{\top}g$ to obtain $y^{(i)}$.
10:     $\tilde{Z}_{i}=SZ_{i}$ and $\tilde{x}^{(i)}\leftarrow\tilde{x}^{(0)}+\tilde{Z}_{1:i}y^{(i)}$.
11:     if the stopping criterion is satisfied then
12:        return  $x=x^{(i)}\leftarrow x^{(0)}+Z_{1:i}y^{(i)}$.
13:     end if
14:     $\tilde{\tau}_{i}\leftarrow\|\tilde{Z}_{1:i}\rVert\|y^{(i)}\rVert/\|\tilde{x}^{(i)}\rVert$
15:     if  $\texttt{tol}_{\tau}\cdot\tilde{\tau}_{i}\geq 1$ then
16:        if $\texttt{tol}_{B}\cdot\kappa(\tilde{Z}_{1:i})\geq 1$ then
17:           $t\leftarrow 2t$
18:        end if
19:        return
20:     end if
21:     if i == m then
22:        $t\leftarrow 2t$
23:     end if
24:  end for
Algorithm 3 One cycle of preconditioned sketched GMRES (sGMRES) with adaptive restarting
6 Numerical Experiments
In this section, we present numerical experiments that demonstrate the results of our stability analysis. All experiments were performed in MATLAB R2023a on a Windows 11 HP laptop with an 11th Gen Intel(R) Core processor with 2.80 GHz and 8 GB of RAM. We used (with modifications) the open source code provided by Güttel and Simunec as part of their recent work [7]. Our code is available at https://github.com/burkel8/stability_sGMRES. The sketching operator used in all experiments is the subsampled random Hadamard transform (see, e.g.,  [2]).
We consider solving a linear system $Ax=b$ with sGMRES using both the truncated Arnoldi process and the sketch-and-select Arnoldi method (sGMRES-trunc and sGMRES-saa, respectively, in the plots). Although a few variants of the sketch-and-select Arnoldi algorithm have been proposed in [7], we use the sGMRES-saa-pinv method, as it was shown to produce the best conditioned basis in their experiments. We also ran the experiments with sGMRES-saa-greedy and obtained similar results, and thus do not include this method here. We note that the different variations of sketch-and-select Arnoldi differ in how the $t$ vectors with which to orthogonalize against are chosen, and we refer to [7] for more details on these methods. In these experiments, we always choose the same value of $t$ in the sketch-and-select Arnoldi algorithm that we use as the truncation parameter in the truncated Arnoldi method. All comparisons to standard GMRES use GMRES with modified Gram-Schmidt. We do not use any preconditioner in the experiments, i.e., $M_{L}=M_{R}=I$. Finally, we drop the $\hat{.}$ notation in this section, as all quantities are computed quantities.
6.1 Illustration of the backward error bound
In this subsection, we use three different examples to illustrate the results of our analysis shown in Sections 4 and 5. For all sketching techniques we take $s=2(m+1)$, where $m$ is the maximum length of each Arnoldi cycle, as suggested in [10]. We display the backward error, $\tau_{i}$ from (43), and the condition numbers of $B_{1:i}$ and $AB_{1:i}$ (approximated using the condition numbers of $SB_{1:i}$ and $SAB_{1:i}$, respectively) at each iteration. These condition number approximations are denoted as $\kappa(B_{1:i})$ and $\kappa(AB_{1:i})$ in the plots.
6.1.1 Attainable worst case for sGMRES
In the first example, the matrix $A$ is a $400\times 400$ random banded matrix with condition number $\kappa(A)=10$, obtained using the MATLAB gallery function with randsvd as input. The right-hand side is taken to be the third singular vector of $A$. We run $m=400$ iterations of standard GMRES and sGMRES with both truncated Arnoldi and sketch-and-select Arnoldi with $t=2$, and do not use any randomized subspace embedding for this example, i.e., $S$ is chosen to be a $400$-by-$400$ identity matrix. Theoretically, the identity matrix is the most stable choice for $S$, but is not an appropriate choice of sketching operator in practice. Here, we use the identity matrix solely to illustrate the worst case for sGMRES, i.e., the case where backward error of sGMRES barely decreases due to the ill-conditioned basis $B_{1:i}$.
The results are shown in Figure 1, where we see that the condition number of $B_{1:i}$ and $\tau_{i}$ generated by the truncated Arnoldi grows to $10^{16}$, leading to a poor backward error result for sGMRES-trunc. In the case of sGMRES-saa, the backward error is approximately $10^{-7}$ as both the condition number of $B_{1:i}$ and $\tau_{i}$ are not too large. For GMRES, it can be seen that the condition number of $B_{1:i}$ remains small, as does $\tau_{i}$, yielding a very small backward error after $400$ iterations. Thus, we can see that the bound in Theorem 1 can be attainable for some cases.
Figure 1: Example 1: From the left to right, the plots show the backward error, $\tau_{i}$ and the approximate condition numbers of basis $B_{i}$ and $AB_{i}$, monitored every iteration, for different implementations of sGMRES, and an implementation of standard GMRES.
6.1.2 Comparison between the bounds shown in Theorem 1 and Lemma 2
In the second example, the matrix $A$ is the Norris/torso3 matrix of size $n=259,156$, and the right-hand side vector is randomly generated according to the standard normal distribution. We take $m=300$ and $t=2$ for this example. We use this example to demonstrate that the bound outlined in Lemma 2 is sharper than the one mentioned in Theorem 1, and proves useful in certain situations.
As shown in Figure 2, the behavior of the backward error for both sGMRES-trunc and sGMRES-saa diverges; notably, while the backward error for sGMRES-saa stops improving after about 160 iterations, sGMRES-trunc continues to reduce its error reaching a small level, despite both methods having ill-conditioned $B_{1:i}$. This indicates that for this scenario, the bounds shown by Theorem 1, specifically (29) and (39), are overly pessimistic and fail to account for the discrepancy observed. We further notice that $\tau_{i}$ for sGMRES-trunc is considerably smaller than for sGMRES-saa, explaining the superior backward error performance of sGMRES-trunc, as anticipated by Lemma 2.
Figure 2: Example 2: From the left to right, the plots show the backward error, $\tau_{i}$ and the approximate condition numbers of the $B_{i}$ and $AB_{i}$, monitored every iteration, for different implementations of sGMRES, and an implementation of standard GMRES.
6.1.3 Comparison between different sGMRES variants with and without restarting
In the third example, the matrix $A$ is the Norris/stomach matrix of size $n=213,360$ obtained from the SuiteSparse matrix collection [4] (also tested in [7]). The right-hand side vector is randomly generated according to the standard normal distribution. We run $\texttt{nrestarts}=5$ cycles of restarted sGMRES with the truncated Arnoldi process, and sketch-and-select Arnoldi, with $m=150$ iterations per cycle, totaling $750$ iterations. We also run restarted GMRES with the same parameters. Additionally, we perform sGMRES with truncated Arnoldi and sketch-and-select Arnoldi with $m=750$ iterations (without restarts). All implementations of sGMRES use a truncation parameter $t=3$.
The results are shown in Figure 3, where we observe that for sGMRES with truncated Arnoldi and sketch-and-select Arnoldi (without restarts), $\tau_{i}$ and the condition numbers of both $B_{1:i}$ and $AB_{1:i}$ grow to about $10^{16}$, causing the backward error to stagnate, in accordance with the results of Theorem 1 and Lemma 2. The restarted implementation helps regain stability, leading to a continued reduction in backward error with each restart, eventually reaching $O(\bm{u})$. This is because restarting can remove the influence of the condition number of $B_{1:i}$ as predicted in Theorem 2.
In Figure 3, we also see that the backward error for sGMRES restarted with a sketch-and-select Arnold reduces faster than for sGMRES restarted with truncated Arnoldi in each cycle. Although the condition numbers of $B_{1:i}$ produced by these two methods eventually become large, the values of $\tau_{i}$ are small. Thus, they can produce small backward errors, which again confirms that the bound in Lemma 2 is sharper than in Theorem 1.
Figure 3: Example 3: From the left to right, the plots show the backward error, $\tau_{i}$, and the approximate condition numbers of $B_{i}$ and $AB_{i}$, monitored every iteration, for different implementations of sGMRES, and an implementation of standard GMRES.
6.2 sGMRES with the adaptively restarting strategy
In the final experiment, we test the various adaptive restarting strategies discussed in Section 3. In particular, we test
  • sGMRES (monitor $\tau_{i}$): monitor $\tau_{i}$, and when it gets too large, restart.
  • sGMRES (monitor $\kappa(B_{1:i})$): monitor the condition number of $B_{1:i}$, and when it gets too large, restart.
  • sGMRES (adapt monitor $\tau_{i}$): monitor $\tau_{i}$, and when it gets too large, double the truncation parameter $t$, and restart. Also restart and double $t$ every $m$ iterations even when $\tau_{i}$ is small.
  • sGMRES (adapt monitor $\kappa(B_{1:i})$): monitor the condition number of $B_{1:i}$, and when it gets too large, double the truncation parameter $t$, and restart. Also restart and double $t$ every $m$ iterations even when the condition number of $B_{1:i}$ is small.
  • sGMRES (adapt monitor $\tau_{i}$ and $\kappa(B_{1:i})$): monitor $\tau_{i}$, and when it gets too large, check the size of $\kappa(B_{1:i})$. If $\kappa(B_{1:i})$ is large, double the truncation parameter $t$, and restart. If $\kappa(B_{1:i})$ is not large enough, just restart without increasing $t$ (this method is shown in Algorithm 3). Also restart and double $t$ every $m$ iterations even when the condition number of $B_{1:i}$ is small.
  • restarted GMRES: restart every $m$ iterations.
We solve a linear system by restarted sGMRES with the truncated Arnoldi process using all the various restarting strategies (beginning with $t=2$ for each method). We then plot the backward error curves produced by sGMRES with each restarting strategy, alongside the backward error curves produced by standard restarted GMRES. In each example, the right-hand side is randomly generated according to the standard normal distribution, and both $\texttt{tol}_{\tau}$ and $\texttt{tol}_{B}$ are taken to be the MATLAB machine precision obtained by the eps command.
For GMRES, we restart a maximum of $d$ times, every $m$ iterations. We then allow a maximum of $md$ iterations for all implementations of sGMRES methods, and each implementation to restart according to its own strategy.
In the first example, the matrix $A$ is the Norris/stomach matrix, $m=150$ and $d=4$. In the second example, the matrix $A$ is the fs_760_1 matrix obtained from SuiteSparse, and $m=30$ and $d=20$. In the final example, the matrix $A$ is taken to be the sherman2 matrix from SuiteSparse, with $m=200$ and $d=5$. The results are shown in Figure 4. The curve for sGMRES (monitor $\tau_{i}$) is hidden behind the curve for sGMRES (adapt monitor $\tau_{i}$) for each of the examples we tested. This is because $\tau_{i}$ is not directly dependent on the condition number of $B_{1:i}$, and so increasing $t$ does not help in this case. From the examples we have tested, sGMRES (adapt monitor $\tau_{i}$ and $\kappa(B_{1:i})$) seems to be the best approach as it produces backward error curves close to standard restarted GMRES in each of the examples.
Figure 4: Adaptive restarting experiment. From left to right, the backward error plots for Norris/Stomach, fs_760_1, and sherman2. The curve for sGMRES (monitor $\tau_{i}$), i.e., the dark blue line, is hidden behind the curve for sGMRES (adapt monitor $\tau_{i}$), i.e., the yellow line, for each of the examples we tested. The curve for sGMRES (adpt monitor $\tau_{i}$ and $\kappa(B_{1:i})$), i.e., the green line, is almost hidden behind the curve for GMRES, i.e., the light blue one.
7 Conclusions
In this paper, we performed a backward stability analysis of sketched GMRES and showed how the backward error is largely influenced by the condition number of the generated Krylov basis. However, in [7, Figure 9], it was demonstrated that the backward error of sGMRES can still decrease even when the basis becomes ill-conditioned. To provide more insight into this finding from [7], we proposed a sharper bound that relies on $\|\hat{B}_{1:k}\rVert\|\hat{y}^{(k)}\rVert/\|\hat{x}^{(k)}\rVert$ rather than $\kappa(\hat{B}_{1:k})$. We then showed how restarting can be used to stabilize a sketched GMRES implementation, and proposed an adaptive restarting strategy for sGMRES based on our analysis. Our numerical experiments indicate that the convergence of sketched GMRES with our adaptive restarting approach often performs comparably to standard GMRES and has the potential to result in substantial time savings.
Acknowledgments
All authors are supported by the Charles University Research Centre program No. UNCE/24/SCI/005 and by the European Union (ERC, inEXASCALE, 101075632). Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.

### v2

@@
 Appl, 2024] for solving linear systems $Ax=b$, and show that the backward stability at iteration $i$ depends on the conditioning of the Krylov basis $B_{1:i}$ as long as the condition number of $AB_{1:i}$ can be bounded by $1/O(\bm{u})$, where $\bm{u}$ is the unit roundoff.
-Under this condition, we show that sketched GMRES is backward stable as long as the condition number of $B_{1:i}$ is not too large.
-Under additional assumptions, we then show that the stability of a restarted implementation of sketched GMRES can be independent of the condition number of $B_{1:i}$, and restarted sketched GMRES is backward stable.
+Under this condition, we demonstrate that the stability of the sketched GMRES is mainly affected by the condition number of $B_{1:i}$, which can be mitigated by using the restarting technique.
 We also derive sharper bounds that explain why the backward error can be small even in cases when the basis $B_{1:i}$ is very ill-conditioned, which has been observed in the literature but not yet explained theoretically.
@@
 1 Introduction
-In this paper, we analyze the numerical stability of sketched GMRES (sGMRES) [10] for solving nonsymmetric and nonsingular linear systems of the form
-$Ax=b,\enspace A\in\mathbb{R}^{n\times n}$ (1)
+In this paper, we analyze the numerical stability of sketched GMRES (sGMRES) [NT2024] for solving nonsymmetric and nonsingular linear systems of the form
+$Ax=b,\kern 5.0ptA\in\mathbb{R}^{n\times n}$ (1)
 with a given matrix $A$ and a right-hand side $b$.
-Although standard GMRES [12] is one of the most popular methods to solve (1), it comes at the expense of constructing a matrix $V_{1:i}\in\mathbb{R}^{n\times i}$ with columns that form an orthonormal basis for the Krylov subspace
+Although standard GMRES [SaadSchultz1986] is one of the most popular methods to solve (1), it comes at the expense of constructing a matrix $V_{1:i}\in\mathbb{R}^{n\times i}$ with columns that form an orthonormal basis for the Krylov subspace
 $\mathcal{K}_{i}(A,r^{(0)})=\operatorname{span}\{r^{(0)},Ar^{(0)},\dotsc,A^{i-1}r^{(0)}\},$
 where $i$ is the iteration number, and $r^{(0)}=b-Ax^{(0)}$ is the initial residual corresponding to an initial approximation $x^{(0)}$.
-Sketched GMRES was proposed in [10] as an attractive alternative to GMRES that exploits the technique of randomized sketching [9] to avoid working with an orthonormal basis.
+Sketched GMRES was proposed in [NT2024] as an attractive alternative to GMRES that exploits the technique of randomized sketching [martinsson2020randomized] to avoid the high cost of constructing an orthonormal basis.
 Instead, the basis is only partially orthogonalized, thereby reducing the arithmetic cost from $O(ni^{2})$ to $O(i^{3}+ni\log i)$.
-It has been observed in numerical experiments in [10] that sGMRES can fail once the condition number of the basis $B_{1:i}$ becomes too large, and methods such as the sketch-and-select Arnoldi algorithm [7] have been proposed as a means to ensure the condition number of the basis $B_{1:i}$ does not grow too rapidly.
-Indeed, numerical experiments in [7] showed that working with a well-conditioned basis improved the overall convergence behavior of sketched GMRES.
-In this paper, we rigorously analyze the backward stability of preconditioned sketched GMRES, and justify the experimental observations made in the literature [10, 7].
+It has been observed in numerical experiments in [NT2024] that sGMRES can fail once the condition number of the basis $B_{1:i}$ becomes too large, and methods such as the sketch-and-select Arnoldi algorithm [guttel2024sketch] have been proposed as a means to ensure the condition number of the basis $B_{1:i}$ does not grow too rapidly.
+Indeed, numerical experiments in [guttel2024sketch] showed that working with a well-conditioned basis improved the overall convergence behavior of sketched GMRES.
+In this paper, we rigorously analyze the backward stability of preconditioned sketched GMRES, and justify the experimental observations made in the literature [NT2024, guttel2024sketch].
 Namely, we show that the backward error of sketched GMRES depends on $\kappa(Z_{1:i})$ as long as $\kappa(AZ_{1:i})$ is not too large, where $Z_{1:i}=M_{R}^{-1}B_{1:i}$ with the right preconditioner $M_{R}$.
@@
 We present numerical experiments to demonstrate the numerical behavior we expect to observe in sketched GMRES as predicted by our new analysis, and demonstrate the improvements in the backward error that can be obtained from the adaptive restart approach.
-[…]
+We briefly outline the notation and some basic definitions used throughout this paper.
+We use MATLAB indexing to denote submatrices.
+For example, we use $X_{1:i}$ to denote the first $i$ columns of $X$.
+We use superscripts to denote iteration numbers.
+For example, $x^{(i)}$ denotes the approximate solution of $Ax=b$ at the $i$-th iteration of GMRES or sketched GMRES.
+We use $\|\cdot\rVert$ to denote the vector $2$ norm and $\bigl\|\cdot\bigr\rVert_{\mathsf{F}}$ to denote the Frobenius norm.
+We use $\bm{u}$ to denote the unit roundoff.
+We also use $\hat{\cdot}$ to denote computed quantities.
+For example, $\hat{x}$ denotes the computed value of the vector $x$ in finite precision.
+The (normwise) backward error of $\hat{x}$ is defined as
 $\min\{\epsilon:(A+\Delta A)\hat{x}=b+\Delta b,\bigl\|\Delta A\bigr\rVert_{\mathsf{F}}\leq\epsilon\bigl\|A\bigr\rVert_{\mathsf{F}},\|\Delta b\rVert\leq\epsilon\|b\rVert\}=\frac{\|b-A\hat{x}\rVert}{\bigl\|A\bigr\rVert_{\mathsf{F}}\|\hat{x}\rVert+\|b\rVert}.$
 The remainder of this paper is organized as follows.
-In Section 2, we give an overview of GMRES and the MOD-GMRES framework [3] for performing a backward stability analysis of different GMRES variants.
+In Section 2, we give an overview of GMRES and the MOD-GMRES framework [BHMV2024] for performing a backward stability analysis of different GMRES variants.
 We then discuss preconditioned sketched GMRES in Section 3, and analyze the backward stability of preconditioned sketched GMRES and restarted sketched GMRES, respectively, in Sections 4 and 5.
@@
 Furthermore, for optimal performance it is common to exploit techniques such as deflation, mixed precision, or randomization within GMRES.
-In order to understand the finite precision behavior of any given GMRES variant, a backward error analysis framework was developed in [3], which simplifies the process of deriving bounds for the attainable normwise backward and forward errors of the computed solutions of different GMRES variants.
+In order to understand the finite precision behavior of any given GMRES variant, a backward error analysis framework was developed in [BHMV2024], which simplifies the process of deriving bounds for the attainable normwise backward and forward errors of the computed solutions of different GMRES variants.
 The framework is built from a backward error analysis of a generic GMRES algorithm known as modular GMRES or MOD-GMRES, that is composed of three elemental operations:
@@
 By specializing these operations and meeting their assumptions, as outlined in the framework, modular GMRES can describe many GMRES implementations and variants, and can be used to perform their backward error analysis.
-For example, well-known results from [5] were derived in [3] using the framework.
+For example, well-known results from [drkovsova1995numerical] were derived in [BHMV2024] using the framework.
 In particular, it was shown that GMRES with Householder orthogonalization (HH-GMRES) produces a computed solution whose backward error is on the order of the unit roundoff, and is thus backward stable.
-Similarly, the backward stability result for GMRES with modified Gram-Schmidt (MGS-GMRES), derived in [11], was also recovered using the framework in [3].
+Similarly, the backward stability result for GMRES with modified Gram-Schmidt (MGS-GMRES), derived in [paige2006modified], was also recovered using the framework in [BHMV2024].
 3 Preconditioned Sketched GMRES
@@
 for $\epsilon\in(0,1)$, and every $v\in\mathcal{K}_{i}(M_{L}^{-1}AM_{R}^{-1},M_{L}^{-1}r^{(0)})$.
-The sketching matrix $S$ needs to be drawn at random to ensure the relation (5) holds with a high probability.
+In practice, the sketching matrix $S$ may not be explicitly available and needs to be drawn at random to achieve (5) with high probability.
+Additionally, $S$ must be constructed without prior knowledge of the Krylov subspace $\mathcal{K}_{i}(M_{L}^{-1}AM_{R}^{-1},M_{L}^{-1}r^{(0)})$ and should typically be a valid embedding for a broad subspace, specifically $\mathbb{R}^{n\times i}$.
 In practice, the sketched least squares problem (4) is computed by solving a triangular linear system
@@
 0: A computed solution $x\in\mathbb{R}^{n}$ approximating the solution of $Ax=b$.
-1:  $r^{(0)}\leftarrow b-Ax^{(0)}$ and $g\leftarrow SM_{L}^{-1}r^{(0)}$.
-2:  $B_{1}\leftarrow r^{(0)}/\|r^{(0)}\rVert$.
+[…]
+[…]
 3:for $i=1:m$ do
@@
 Algorithm 2 Restarted preconditioned sketched GMRES
-It is clear from experimental observations made in [10, 7] that the basis condition number should not grow too large throughout the iteration in order to ensure numerical stability of sketched GMRES.
-The non-orthogonal basis can be constructed using the $t$ truncated Arnoldi process, where each new Krylov vector is orthogonalized against the previous $t$ vectors.
-Alternatively, the sketch-and-select method [7] uses the sketched basis $SB_{i}$ to select $t$ vectors, among all previously computed vectors, with which to perform the partial orthogonalization, with the aim of generating a basis with smaller condition number than that generated by truncated Arnoldi.
+It is clear from experimental observations made in [NT2024, guttel2024sketch] that the basis condition number should not grow too large throughout the iteration in order to ensure numerical stability of sketched GMRES.
+The non-orthogonal basis can be constructed using the $t$ truncated Arnoldi process, as summarized in Algorithm 3, where each new Krylov vector is orthogonalized against the previous $t$ vectors.
+Alternatively, the sketch-and-select method [guttel2024sketch] uses the sketched basis $SB_{i}$ to select $t$ vectors, among all previously computed vectors, with which to perform the partial orthogonalization, with the aim of generating a basis with smaller condition number than that generated by truncated Arnoldi.
+0: A matrix $A\in\mathbb{R}^{n\times n}$, a truncated parameter $t$, the basis $B_{1:i-1}$ generated by the first i1i-1 truncated Arnoldi iterations.
+0: The basis $B_{1:i}$.
+[…]
+2:for $l=\max(1,i-t+1):i$ do
+3:  $B_{i}=B_{i}-B_{l}(B_{l}^{\top}B_{i})$.
+4:end for
+[…]
+Algorithm 3 The $i$-th iteration of the truncated Arnoldi process
 4 Backward stability of preconditioned sketched GMRES
-In this section, we aim to analyze the backward stability of preconditioned sketched GMRES (Algorithm 1) by primarily following the analysis of the MOD-GMRES framework [3].
+In this section, we aim to analyze the backward stability of preconditioned sketched GMRES (Algorithm 1) by primarily following the analysis of the MOD-GMRES framework [BHMV2024].
 The first $i$ iterations of preconditioned sketched GMRES can be summarized in the following steps:
@@
 $\hat{W}_{1:i}=M_{L}^{-1}A\hat{Z}_{1:i}+\Delta W_{1:i},\quad\bigl\|\Delta W_{1:i}\bigr\rVert_{\mathsf{F}}\leq\varepsilon_{W}\|M_{L}^{-1}\rVert\bigl\|A\bigr\rVert_{\mathsf{F}}\bigl\|\hat{Z}_{1:i}\bigr\rVert_{\mathsf{F}},$ (7)
-with $\varepsilon_{W}\leq O(\bm{u})$ from the standard rounding error analysis of [8].
+with $\varepsilon_{W}\leq O(\bm{u})$ from the standard rounding error analysis of [H2002].
 Similarly, we assume that the computed result of $\tilde{b}:=M_{L}^{-1}b$ computed by Line 1 in Algorithm 1 satisfies
@@
 Step 2: Least squares solver
-In iteratively solving the least squares problem in Step 2, we assume that there is a special dimension $k\leq n$, known as the key dimension [3], for which the computed solution is guaranteed to have reached a small backward error, i.e., $\hat{b}$ lies in the range of non-singular basis $\hat{W}_{1:k}$.
+In iteratively solving the least squares problem in Step 2, we assume that there is a special dimension $k\leq n$, known as the key dimension [BHMV2024], for which the computed solution is guaranteed to have reached a small backward error, i.e., $\hat{b}$ lies in the range of non-singular basis $\hat{W}_{1:k}$.
 The definition of the key dimension is as follows:
-$\sigma_{\min}(\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix})$ $\leq\omega_{\kappa(r,W)}\bigl{\|}\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix}\bigr{\rVert}_{\mathsf{F}},$ (9)
+[…]
 $\sigma_{\min}(\hat{W}_{1:k})$ $\geq\omega_{\kappa(W)}\bigl\|\hat{W}_{1:k}\bigr\rVert_{\mathsf{F}}$ (10)
@@
 We then provide a bound on this sketched residual in Lemma 1, and use this bound in the proof of our final backward stability result in Theorem 2.
-We first set up the least squares problem by constructing the matrix $\hat{C}_{1:k}$, and the vector $g:=SM_{L}^{-1}b$. We assume that
-$\hat{C}_{1:k}$ $=S\hat{W}_{1:k}+\Delta C_{1:k},\quad\|\Delta C_{j}\rVert\leq\varepsilon_{S}\|S\hat{W}_{j}\rVert,\quad\forall j\leq k,$ (11)
+We first set up the least squares problem by constructing the matrix $\hat{C}_{1:i}$, and the vector $g:=SM_{L}^{-1}b$. We assume that
+$\hat{C}_{1:i}$ $=S\hat{W}_{1:i}+\Delta C_{1:i},\quad\|\Delta C_{j}\rVert\leq\varepsilon_{S}\|S\hat{W}_{j}\rVert,\quad\forall j\leq i,$ (11)
 $\hat{g}$ $=S\hat{b}+\Delta g,\quad\|\Delta g\rVert\leq\varepsilon_{S}\|M_{L}^{-1}\rVert\|S\hat{b}\rVert$, (12)
-and furthermore, that the computed solution $\hat{y}^{(k)}$ of the least squares problem in Step 2 satisfies
-$\begin{split}&\hat{y}^{(k)}=\operatorname*{\arg\min}_{y}\|\hat{g}+\Delta\hat{g}-(\hat{C}_{1:k}+\Delta\hat{C}_{1:k})y\rVert,\\ &\|\begin{bmatrix}\Delta\hat{g}&\Delta\hat{C}_{1:k}\end{bmatrix}e_{j}\rVert\leq\varepsilon_{sls}\|\begin{bmatrix}\hat{g}&\hat{C}_{1:k}\end{bmatrix}e_{j}\rVert,\quad\forall j\leq k+1,\end{split}$ (13)
-with $\varepsilon_{sls}\leq O(\bm{u})$ from [8, Theorem 20.3]. Substituting $\hat{C}_{1:k}$ and $\hat{g}$ in (13) with (11) and (12), we have
-$\begin{split}&\hat{y}^{(k)}=\operatorname*{\arg\min}_{y}\big{\|}S\tilde{\hat{b}}+\Delta\tilde{g}-(S\hat{W}_{1:k}+\Delta\tilde{C}_{1:k})y\big{\rVert},\\ &\|\Delta\tilde{g}\rVert\leq\|\Delta g\rVert+\|\Delta\hat{g}\rVert\leq(1+\varepsilon_{sls})\varepsilon_{S}\|S\hat{b}\rVert+\varepsilon_{sls}\|S\hat{b}\rVert\leq\bigl{(}\varepsilon_{sls}+(1+\varepsilon_{sls})\varepsilon_{S}\bigr{)}\|S\hat{b}\rVert,\\ &\|\Delta\tilde{C}_{j}\rVert\leq\|\Delta C_{j}\rVert+\|\Delta\hat{C}_{j}\rVert\leq\bigl{(}\varepsilon_{sls}+(1+\varepsilon_{sls})\varepsilon_{S}\bigr{)}\|S\hat{W}_{j}\rVert,\quad\forall j\leq k.\end{split}$ (14)
-We can now prove a bound on the norm of the sketched residual at iteration $k$, i.e., the norm of the vector
+and furthermore, that the computed solution $\hat{y}^{(i)}$ of the least squares problem in Step 2 satisfies
+$\begin{split}&\hat{y}^{(i)}=\operatorname*{\arg\min}_{y}\|\hat{g}+\Delta\hat{g}-(\hat{C}_{1:i}+\Delta\hat{C}_{1:i})y\rVert,\\ &\|\begin{bmatrix}\Delta\hat{g}&\Delta\hat{C}_{1:i}\end{bmatrix}e_{j}\rVert\leq\varepsilon_{sls}\|\begin{bmatrix}\hat{g}&\hat{C}_{1:i}\end{bmatrix}e_{j}\rVert,\quad\forall j\leq i+1,\end{split}$ (13)
+with $\varepsilon_{sls}\leq O(\bm{u})$ from [H2002, Theorem 20.3]. Substituting $\hat{C}_{1:i}$ and $\hat{g}$ in (13) with (11) and (12), we have
+$\begin{split}&\hat{y}^{(i)}=\operatorname*{\arg\min}_{y}\big\|S\tilde{\hat{b}}+\Delta\tilde{g}-(S\hat{W}_{1:i}+\Delta\tilde{C}_{1:i})y\big\rVert,\\ &\|\Delta\tilde{g}\rVert\leq\|\Delta g\rVert+\|\Delta\hat{g}\rVert\leq(1+\varepsilon_{sls})\varepsilon_{S}\|S\hat{b}\rVert+\varepsilon_{sls}\|S\hat{b}\rVert\leq\bigl(\varepsilon_{sls}+(1+\varepsilon_{sls})\varepsilon_{S}\bigr)\|S\hat{b}\rVert,\\ &\|\Delta\tilde{C}_{j}\rVert\leq\|\Delta C_{j}\rVert+\|\Delta\hat{C}_{j}\rVert\leq\bigl(\varepsilon_{sls}+(1+\varepsilon_{sls})\varepsilon_{S}\bigr)\|S\hat{W}_{j}\rVert,\quad\forall j\leq i.\end{split}$ (14)
+We can now prove a bound on the norm of the sketched residual at the key dimension $k$, i.e., the norm of the vector
 $r^{(k)}_{S}:=S\hat{b}+\Delta\tilde{g}-(S\hat{W}_{1:k}+\Delta\tilde{C}_{1:k})\hat{y}^{(k)}.$ (15)
@@
 Assume that $\hat{W}_{1:k}$ satisfies
-$\sigma_{\min}(\begin{bmatrix}\phi S\hat{b}&S\hat{W}_{1:k}\end{bmatrix})$ $\leq c_{1}(n,k)\bigl{(}\varepsilon_{S}+\varepsilon_{sls}\bigr{)}\bigl{\|}\begin{bmatrix}\phi S\hat{b}&S\hat{W}_{1:k}\end{bmatrix}\bigr{\rVert}_{\mathsf{F}},$ (16)
+[…]
 $\sigma_{\min}(S\hat{W}_{1:k})$ $\geq c_{2}(n,k)\bigl(\varepsilon_{S}+\varepsilon_{sls}\bigr)\bigl\|S\hat{W}_{1:k}\bigr\rVert_{\mathsf{F}},$ (17)
@@
 Proof.
-We follow the proof in [3] and employ [11, Theorem 2.4] to bound $\|r^{(k)}_{S}\rVert$. Applying [11, Theorem 2.4] to the least squares problem (14), we have
+We follow the proof in [BHMV2024] and employ [paige2006modified, Theorem 2.4] to bound $\|r^{(k)}_{S}\rVert$. Applying [paige2006modified, Theorem 2.4] to the least squares problem (14), we have
 $\|r^{(k)}_{S}\rVert^{2}\leq\sigma_{\min}^{2}\bigl(\begin{bmatrix}\phi(S\hat{b}+\Delta\tilde{g})&S\hat{W}_{1:k}+\Delta\tilde{C}_{1:k}\end{bmatrix}\bigr)\biggl(\frac{1}{\phi^{2}}+\frac{\|\hat{y}^{(k)}\rVert^{2}}{1-\delta_{k}^{2}}\biggr),$ (19)
@@
 $\delta_{k}=\frac{\sigma_{\min}\bigl(\begin{bmatrix}\phi(S\hat{b}+\Delta\tilde{g})&S\hat{W}_{1:k}+\Delta\tilde{C}_{1:k}\end{bmatrix}\bigr)}{\sigma_{\min}\bigl(S\hat{W}_{1:k}+\Delta\tilde{C}_{1:k}\bigr)}<1.$ (20)
-We choose $\phi$ such that $\phi^{-2}=\|\hat{y}^{(k)}\rVert^{2}/(1-\delta_{k}^{2})$ as in [3, (3.17)], which can simplify (19) as
+We choose $\phi$ such that $\phi^{-2}=\|\hat{y}^{(k)}\rVert^{2}/(1-\delta_{k}^{2})$ as in [BHMV2024, (3.17)], which can simplify (19) as
 $\|r^{(k)}_{S}\rVert\leq\sqrt{2}\,\sigma_{\min}\bigl(\begin{bmatrix}\phi(S\hat{b}+\Delta\tilde{g})&S\hat{W}_{1:k}+\Delta\tilde{C}_{1:k}\end{bmatrix}\bigr)\phi^{-1}.$ (21)
@@
 By continuity, this implies that there exists $\phi\in(0,\|\hat{y}^{(k)}\rVert^{-1})$ such that $\phi^{-2}=\|\hat{y}^{(k)}\rVert^{2}/(1-\delta_{k}^{2})$ and $\delta_{k}<1$.
-It remains to bound $\sigma_{\min}\bigl{(}\begin{bmatrix}\phi(S\hat{b}+\Delta\tilde{g})&S\hat{W}_{1:k}+\Delta\tilde{C}_{1:k}\end{bmatrix}\bigr{)}$ and $\phi$ in (21).
-By the assumption (16), and the estimation of $\|\Delta\tilde{g}\rVert$ and $\|\Delta\tilde{C}_{1:k}\rVert$ shown in (14), we have
+[…]
 $\begin{split}\sigma_{\min}\bigl(&\begin{bmatrix}\phi(S\hat{b}+\Delta\tilde{g})&S\hat{W}_{1:k}+\Delta\tilde{C}_{1:k}\end{bmatrix}\bigr)\\ &\leq c_{1}(n,k)\bigl(\varepsilon_{S}+\varepsilon_{sls}\bigr)\bigl\|\begin{bmatrix}\phi S\hat{b}&S\hat{W}_{1:k}\end{bmatrix}\bigr\rVert_{\mathsf{F}}\\ &\leq c_{1}(n,k)\bigl(\varepsilon_{S}+\varepsilon_{sls}\bigr)\bigl(\|S\hat{b}\rVert\phi+\bigl\|S\hat{W}_{1:k}\bigr\rVert_{\mathsf{F}}\bigr).\end{split}$ (23)
@@
 Step 3: Computation of the approximate solution
-For $\hat{Z}_{1:k}\in\mathbb{R}^{n\times k}$ and $\hat{y}^{(k)}\in\mathbb{R}^{k}$, we assume that
-$\hat{x}^{(k)}=\hat{Z}_{1:k}\hat{y}^{(k)}+\Delta x^{(k)},\quad\|\Delta x^{(k)}\rVert\leq\varepsilon_{x}\|\hat{Z}_{1:k}\rVert\|\hat{y}^{(k)}\rVert$ (27)
-with $\varepsilon_{x}\leq O(\bm{u})$ from [8].
+For $\hat{Z}_{1:i}\in\mathbb{R}^{n\times i}$ and $\hat{y}^{(i)}\in\mathbb{R}^{i}$, we assume that
+$\hat{x}^{(i)}=\hat{Z}_{1:i}\hat{y}^{(i)}+\Delta x^{(i)},\quad\|\Delta x^{(i)}\rVert\leq\varepsilon_{x}\|\hat{Z}_{1:i}\rVert\|\hat{y}^{(i)}\rVert$ (27)
+with $\varepsilon_{x}\leq O(\bm{u})$ from [H2002].
 We are now ready to describe the backward stability of preconditioned sketched GMRES (Algorithm 1) in the following theorem.
 Theorem 1.
-Assume that Algorithm 1 is applied with a computed preconditioned basis $\hat{Z}_{1:k}\in\mathbb{R}^{n\times k}$ and an $\epsilon$-subspace embedding $S\in\mathbb{R}^{s\times n}$, where (7), (8), (11), (12), (13), and (27) are satisfied.
+Assume that Algorithm 1 is applied with a computed preconditioned basis $\hat{Z}_{1:i}\in\mathbb{R}^{n\times i}$ and an $\epsilon$-subspace embedding $S\in\mathbb{R}^{s\times n}$ satisfying (5) with high probability, where (7), (8), (11), (12), (13), and (27) are satisfied.
 Then there exists $k\leq n$ such that $c_{1}(n,k)\bigl(\varepsilon_{S}+\varepsilon_{sls}\bigr)\sqrt{\frac{1-\epsilon}{1+\epsilon}}\kappa(\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix})\geq 1$.
@@
 Proof.
-The residual of $\hat{x}^{(k)}$ can be bounded by the preconditioned residual as follows:
+Note that $\hat{b}\in\mathbb{R}^{n}$, and we can establish the existence of $k$, since the condition $c_{1}(n,k)\bigl(\varepsilon_{S}+\varepsilon_{sls}\bigr)\sqrt{\frac{1-\epsilon}{1+\epsilon}}\kappa(\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix})\geq 1$ is satisfied when $k=n$.
+Then we only need to prove (29). The residual of $\hat{x}^{(k)}$ can be bounded by the preconditioned residual as follows:
 $\|b-A\hat{x}^{(k)}\rVert=\|M_{L}\bigl(M_{L}^{-1}b-M_{L}^{-1}A\hat{x}^{(k)})\bigr\rVert\leq\|M_{L}\rVert\|M_{L}^{-1}b-M_{L}^{-1}A\hat{x}^{(k)}\rVert,$ (31)
@@
 $\|M_{L}^{-1}b-M_{L}^{-1}A\hat{x}^{(k)}\rVert\leq\|\hat{b}-\hat{W}_{1:k}\hat{y}^{(k)}\rVert+\|\Delta W_{1:k}\hat{y}^{(k)}\rVert+\|\Delta b\rVert+\|M_{L}^{-1}\rVert\|A\Delta x^{(k)}\rVert.$ (32)
-Furthermore, from [10, Fact 2.2] and the definition (15) of $r^{(k)}_{S}$, we have
+Furthermore, from [NT2024, Fact 2.2] and the definition (15) of $r^{(k)}_{S}$, we have
 $\begin{split}\|\hat{b}-\hat{W}_{1:k}\hat{y}^{(k)}\rVert&\leq\frac{1}{1-\epsilon}\|S\hat{b}-S\hat{W}_{1:k}\hat{y}^{(k)}\rVert\\ &\leq\frac{1}{1-\epsilon}\|r^{(k)}_{S}\rVert+\frac{1}{1-\epsilon}\bigl(\|\Delta\tilde{g}\rVert+\bigl\|\Delta\tilde{C}_{1:k}\bigr\rVert_{\mathsf{F}}\|\hat{y}^{(k)}\rVert\bigr).\end{split}$ (33)
@@
 To apply Lemma 1 to bound $\|r^{(k)}_{S}\rVert$, we first need to verify (16) and (17).
-From [1, Corollary 2.2], the condition numbers of the sketched matrices satisfy
+From [BG2022, Corollary 2.2], the condition numbers of the sketched matrices satisfy
 $\kappa(S\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix})\geq\sqrt{\frac{1-\epsilon}{1+\epsilon}}\kappa(\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix})\quad\text{and}\quad\kappa(S\hat{W}_{1:k})\leq\sqrt{\frac{1+\epsilon}{1-\epsilon}}\kappa(\hat{W}_{1:k}),$
 which implies, from $c_{1}(n,k)\bigl(\varepsilon_{S}+\varepsilon_{sls}\bigr)\sqrt{\frac{1-\epsilon}{1+\epsilon}}\kappa(\begin{bmatrix}\phi b&\hat{W}_{1:k}\end{bmatrix})\geq 1$ and (28),
-$\sigma_{\min}(S\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix})$ $\leq c_{1}(n,k)\bigl{(}\varepsilon_{S}+\varepsilon_{sls}\bigr{)}\|S\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix}\rVert$
+[…]
 $\leq c_{1}(n,k)\bigl(\varepsilon_{S}+\varepsilon_{sls}\bigr)\bigl\|S\begin{bmatrix}\phi b&\hat{W}_{1:k}\end{bmatrix}\bigr\rVert_{\mathsf{F}},$
@@
 Remark 1.
-If the preconditioners satisfy $M_{L}=M_{R}=I$, then the relative backward error can be bounded as
+Theorem 1 indicates that (29) and (30) are satisfied with high probability.
+This is due to the $\epsilon$-subspace embedding, which generally meets the condition (5) with high probability.
+Practically, the sketching method rarely fails, and slightly increasing the embedding dimension can further improve its reliability.
+In the worst case, $\epsilon$ may approach $1$ and the constant $(1+\epsilon)/(1-\epsilon)$ may be relatively large.
+For this reason, it is distracting to quantify the failure probabilities; see details in [MT2020].
+Remark 2.
+Let $\varepsilon_{*}$ denote all the $\varepsilon$ with any subscript mentioned in this section.
+Assuming $\varepsilon_{*}\leq O(\bm{u})$, we simplify Theorem 1 for the case without preconditioning, i.e., $M_{L}=M_{R}=I$ and $\kappa(\hat{Z}_{1:k})=\kappa(\hat{B}_{1:k})$: if it holds for the key dimension $k$ that
+$O(\bm{u})\sqrt{\frac{1+\epsilon}{1-\epsilon}}\kappa(\hat{W}_{1:k})\leq 1/2,\qquad O(\bm{u})\kappa(\hat{B}_{1:k})\leq 1/2,$ (39)
+then the relative backward error can be bounded as
 $\frac{\|b-A\hat{x}^{(k)}\rVert}{\bigl\|A\bigr\rVert_{\mathsf{F}}\|\hat{x}^{(k)}\rVert+\|b\rVert}\leq\frac{1+\epsilon}{1-\epsilon}O(\bm{u})\kappa(\hat{B}_{1:k}).$ (40)
 Remark 3.
-Let $\varepsilon_{*}$ denote all the $\varepsilon$ with any subscript mentioned in this section.
 Considering $\varepsilon_{*}\leq O(\bm{u})$, note that the assumption (28) in Theorem 1 can be replaced by
@@
 from (7).
-Therefore, if $S$ is chosen to be a $n\times n$ identity matrix and $\hat{B}_{1:k}$ is chosen to be an orthonormal Krylov basis, Theorem 1 recovers a similar backward stability result as [3, Theorem 3.1].
+Therefore, if $S$ is chosen to be a $n\times n$ identity matrix and $\hat{B}_{1:k}$ is chosen to be an orthonormal Krylov basis, Theorem 1 recovers a similar backward stability result as [BHMV2024, Theorem 3.1].
 Notice that $\kappa(\hat{W}_{1:k})\leq 3\kappa(M_{L}^{-1}AZ_{1:j})$ from (41) if $O(\bm{u})\kappa(M_{L})\kappa(A)\kappa(\hat{Z}_{1:k})\leq 1$.
@@
 In particular, sGMRES is backward stable provided $\kappa(Z_{1:j})$ is not too large.
-In has been observed in numerical experiments in [6, Figure 9] that in some cases, sGMRES will continue to converge even when $\kappa(\hat{Z}_{1:k})>10^{15}$, and so the bound shown in Theorem 1 is not tight.
+It has been observed in numerical experiments in [guttel2024sketch, Figure 9] that in some cases, sGMRES will continue to converge even when $\kappa(\hat{Z}_{1:k})>10^{15}$, and so the bound shown in Theorem 1 is not tight.
 Therefore, we present the following lemma to show a tighter bound, which can be directly derived from (37) by noticing
@@
 $\frac{\|b-A\hat{x}^{(k)}\rVert}{\bigl\|A\bigr\rVert_{\mathsf{F}}\|\hat{x}^{(k)}\rVert+\|b\rVert}\leq\frac{1+\epsilon}{1-\epsilon}O(\bm{u})\frac{\|\hat{B}_{1:k}\rVert\|\hat{y}^{(k)}\rVert}{\|\hat{x}^{(k)}\rVert}.$ (44)
-5 Adaptively restarting sGMRES
-Similar to the case of standard GMRES, restarting can be regarded as an iterative refinement process, and can remove the influence of both $\kappa(M_{L})$ and $\kappa(Z_{1:j})$ on the backward stability of sGMRES, as shown in the following version of [3, Theorem 4.1] describing the stability of restarted sketched GMRES (Algorithm 2).
+5 Adaptive restarting sGMRES
+Similar to the case of standard GMRES, restarting can be regarded as an iterative refinement process as follows:
+  1. 1.
+    Compute the residual $r^{(i)}=b-Ax^{(i)}$.
+  2. 2.
+    Solve $Ad^{(i)}=r^{(i)}$, which is computed by Algorithm 1.
+  3. 3.
+    Update the solution $x^{(i+1)}=x^{(i)}+d^{(i)}$.
+Furthermore, it can remove the influence of both $\kappa(M_{L})$ and $\kappa(Z_{1:j})$ on the backward stability of sGMRES, as shown in the following version of [BHMV2024, Theorem 4.1] describing the stability of restarted sketched GMRES (Algorithm 2).
 Theorem 2.
-Assume that Algorithm 1 is restarted, such that for restart number $i\geq 1$, we compute a basis $\hat{Z}_{1:k^{(i)}}^{(i)}\in\mathbb{R}^{n\times k^{(i)}}$ and a randomized subspace embedding $S^{(i)}\in\mathbb{R}^{s\times n}$, where (7), (8), (11), (12), (13), and (27) are satisfied for $\varepsilon_{W}^{(i)}$, $\varepsilon_{sls}^{(i)}$, $\varepsilon_{S}^{(i)}$, $\varepsilon_{x}^{(i)}\leq O(\bm{u})$.
+Assume that Algorithm 1 is restarted, such that for restart number $i\geq 1$, we compute a basis $\hat{Z}_{1:k^{(i)}}^{(i)}\in\mathbb{R}^{n\times k^{(i)}}$ and a randomized subspace embedding $S\in\mathbb{R}^{s\times n}$ satisfying (5) with high probability, where (7), (8), (11), (12), (13), and (27) are satisfied for $\varepsilon_{W}^{(i)}$, $\varepsilon_{sls}^{(i)}$, $\varepsilon_{S}^{(i)}$, $\varepsilon_{x}^{(i)}\leq O(\bm{u})$.
 Then there exists $k^{(i)}\leq n$ such that
-$O(\bm{u})\sqrt{\frac{1-\epsilon}{1+\epsilon}}\kappa\bigl{(}\begin{bmatrix}\phi r^{(i)}&\hat{W}_{1:k^{(i)}}^{(i)}\end{bmatrix}\bigr{)}\geq 1.$
+$O(\bm{u})\sqrt{\frac{1-\epsilon}{1+\epsilon}}\kappa\bigl(\begin{bmatrix}\phi r^{(i)}&\hat{W}_{1:k^{(i)}}^{(i)}\end{bmatrix}\bigr)\geq 1.$ (45)
 Assume that
-$\Lambda^{(i)}_{1}=O(\bm{u})\frac{1+\epsilon}{1-\epsilon}\kappa(\hat{Z}_{1:k^{(i)}}^{(i)})\kappa_{\mathsf{F}}(A)\ll 1$
-holds for $i$. Then, with high probability, the backward error is reduced in the $i$-th restart cycle by a factor $\Lambda^{(i)}_{1}$ until it holds that
+$\Lambda^{(i)}_{1}=\frac{1+\epsilon}{1-\epsilon}O(\bm{u})\kappa(M_{L})\bigl(1+\bar{\tau}_{k^{(i)}}\bigr)\ll 1$ (46)
+holds for $i$ with
+$\bar{\tau}_{k^{(i)}}=\frac{\|\hat{Z}^{(i)}_{1:k^{(i)}}\rVert\bigl\|A\bigr\rVert_{\mathsf{F}}\|\hat{y}^{(k^{(i)})}\rVert}{\|A\hat{d}^{(i)}\rVert}.$
+Then, with high probability, the backward error is reduced in the $i$-th restart cycle by a factor $\Lambda^{(i)}_{1}$ until it holds that
 $\frac{\|b-A\hat{x}\rVert}{\bigl\|A\bigr\rVert_{\mathsf{F}}\|\hat{x}\rVert+\|b\rVert}\leq O(\bm{u}).$ (47)
-As a consequence of Theorem 2, it is clear that a restarted variant of sketched GMRES (Algorithm 2) may allow us to recover the stability of sGMRES when the conditioning of the basis grows large.
-It is clear from our analysis that care should be taken to ensure $\kappa(Z_{1:i})$ does not get too large as the iteration progresses.
-In fact, as noted in [10], due to the subspace embedding property (5), the condition number of $Z_{1:i}$ can be monitored inexpensively by explicitly computing the condition number of $SZ_{1:i}$.
-The numerical stability of sGMRES can thus be controlled by monitoring the condition number of $SZ_{1:i}$ in every iteration, and restarting when it grows above some specified tolerance $\texttt{tol}_{B}$.
-For the case of a truncated Arnoldi implementation, or the sketch-and-select Arnoldi implementations, the truncation parameter $t$ can then be increased upon restart to ensure the condition number of the basis $Z_{1:i}$ does not grow as rapidly as it did in the previous cycle.
-As an alternative, due to the result of Lemma 2, we may also opt to monitor
-$\tau_{i}:=\frac{\|Z_{1:i}\rVert\|y^{(i)}\rVert}{\|x^{(i)}\rVert},$
-from (43) at each iteration, since it provides a sharper bound on the backward error, and can be inexpensively estimated as
-$\tilde{\tau}_{i}:=\frac{\|SZ_{1:i}\rVert\|y^{(i)}\rVert}{\|Sx^{(i)}\rVert}.$
-We may then opt to restart when $\texttt{tol}_{\tau}\cdot\tau_{i}\geq 1$, for some chosen tolerance $\texttt{tol}_{\tau}$.
-Again, optionally, we may then wish to increase the value of the truncation parameter $t$.
-In Algorithm 3, we present a hybrid approach, where we monitor $\tilde{\tau}_{i}$, and if it is large, we then restart.
-Before restarting, we also check if $SB_{1:i}$ is large, and if it is, we increase the truncation parameter, by, e.g., doubling it.
+Proof.
+Similar to the proof of [BHMV2024, Theorem 4.1], we will employ [BHMV2024, Lemma 4.2], which requires verifying the assumption [BHMV2024, (4.14)], i.e., bounding $\|\hat{r}^{(i)}-A\hat{d}^{(i)}\rVert$ by $\|b-A\hat{x}^{(i)}\rVert$ and $\|b\rVert+\bigl\|A\bigr\rVert_{\mathsf{F}}\|\hat{x}^{(i)}\rVert$.
+From Lemma 2, we have
+$\begin{split}\|\hat{r}^{(i)}-A\hat{d}^{(i)}\rVert&\leq\frac{1+\epsilon}{1-\epsilon}O(\bm{u})\kappa(M_{L})\biggl(\frac{\|\hat{Z}^{(i)}_{1:k^{(i)}}\rVert\bigl\|A\bigr\rVert_{\mathsf{F}}\|\hat{y}^{(k^{(i)})}\rVert}{\|A\hat{d}^{(i)}\rVert}\|A\hat{d}^{(i)}\rVert+\|\hat{r}^{(i)}\rVert\biggr)\\ &\leq\frac{1+\epsilon}{1-\epsilon}O(\bm{u})\kappa(M_{L})\biggl(\bar{\tau}_{k^{(i)}}\bigl(\|\hat{r}^{(i)}-A\hat{d}^{(i)}\rVert+\|\hat{r}^{(i)}\rVert\bigr)+\|\hat{r}^{(i)}\rVert\biggr)\\ &\leq\frac{1+\epsilon}{1-\epsilon}O(\bm{u})\kappa(M_{L})\biggl(\bigl(1+\bar{\tau}_{k^{(i)}}\bigr)\|\hat{r}^{(i)}\rVert+\bar{\tau}_{k^{(i)}}\|\hat{r}^{(i)}-A\hat{d}^{(i)}\rVert\biggr).\end{split}$ (48)
+By moving the term related to $\|\hat{r}^{(i)}-A\hat{d}^{(i)}\rVert$ from the right-hand side of (48) to the left-hand side and using assumption (46), we further obtain
+$\begin{split}\|\hat{r}^{(i)}-A\hat{d}^{(i)}\rVert&\leq\frac{1+\epsilon}{1-\epsilon}O(\bm{u})\kappa(M_{L})\bigl(1+\bar{\tau}_{k^{(i)}}\bigr)\|\hat{r}^{(i)}\rVert\\ &\leq\frac{1+\epsilon}{1-\epsilon}O(\bm{u})\kappa(M_{L})\bigl(1+\bar{\tau}_{k^{(i)}}\bigr)\bigl(\|b-A\hat{x}^{(i)}\rVert\\ &\quad+O(\bm{u})(\|b\rVert+\bigl\|A\bigr\rVert_{\mathsf{F}}\|\hat{x}^{(i)}\rVert)\bigr),\end{split}$ (49)
+which gives [BHMV2024, (4.14)]. Together with [BHMV2024, Lemma 4.2], we can draw the conclusion. ∎
+Remark 6.
+When no preconditioning is used, i.e., $M_{L}=M_{R}=I$ and $\hat{Z}_{1:k^{(i)}}^{(i)}=\hat{B}_{1:k^{(i)}}^{(i)}$, if it holds for the key dimension $k^{(i)}$ that
+$\Lambda^{(i)}_{1}=O(\bm{u})\frac{1+\epsilon}{1-\epsilon}\biggl(1+\frac{\|\hat{B}^{(i)}_{1:k^{(i)}}\rVert\bigl\|A\bigr\rVert_{\mathsf{F}}\|\hat{y}^{(k^{(i)})}\rVert}{\|A\hat{d}^{(i)}\rVert}\biggr)\ll 1,$ (50)
+then, with high probability, the backward error is reduced in the $i$-th restart cycle by a factor $\Lambda^{(i)}_{1}$ until it holds that
+$\frac{\|b-A\hat{x}\rVert}{\bigl\|A\bigr\rVert_{\mathsf{F}}\|\hat{x}\rVert+\|b\rVert}\leq O(\bm{u}).$ (51)
+Remark 7.
+As a consequence of Theorem 2, it is clear that a restarted variant of sketched GMRES (Algorithm 2) may allow us to recover the stability of sGMRES when $\bar{\tau}_{k^{(i)}}$ is moderate.
+However, it should be noted that the assumptions, particularly (45), of this theorem are overly pessimistic.
+In general, selecting a sufficiently large $m$ is required to guarantee (45), but, in practice, we typically select a relatively small value for the maximum iterations per restart cycle, which makes it difficult to satisfy (45).
+As discussed in Remark 7, for restarted sGMRES, there are two possible reasons for instability: One is due to choosing an $m$ that is too small; the other is due to using a value of $t$ that is too small, which has an influence on $\bar{\tau}_{k^{(i)}}$.
+Since the choice of $m$ depends mainly on storage in practice, here we only consider recovering from the instability caused by a small $t$.
+Employing a small $t$ can result in considerable acceleration, though it might also introduce instability issues.
+Next, our aim is to propose an adaptive restarting strategy to recover from the instability of sGMRES caused by using a small $t$.
+The definition of $\Lambda_{1}^{(1)}$ in Theorem 2 relying on $\bar{\tau}_{k^{(i)}}$, which influences the reduced factor of the backward error in the $i$-th restart cycle, implies that care should be taken to ensure $\bar{\tau}_{k^{(i)}}$ does not get too large as the iteration progresses.
+In fact, as noted in [NT2024], due to the subspace embedding property (5), $\bar{\tau}_{i}$ can be inexpensively estimated as
+$\tilde{\tau}_{i}:=\frac{\|SZ_{1:i}\rVert\bigl\|A\bigr\rVert_{\mathsf{F}}\|y^{(i)}\rVert}{\|SAZ_{1:i}y^{(i)}\rVert}.$
+The numerical stability of sGMRES can thus be controlled by monitoring $\tilde{\tau}_{i}$ at each iteration, and doubling the value of $t$ when it grows above some specified tolerance.
+We then proceed to double $t$ again only if $\tilde{\tau}_{i}$ is still large and $\tilde{\tau}_{i}$ significantly exceeds $\tilde{\tau}_{i-1}$, indicating that $t$ remains insufficiently large.
+Restarted sGMRES combined with this adaptive strategy is presented in Algorithm 4.
 0: A matrix $A\in\mathbb{R}^{n\times n}$, a right-hand side $b\in\mathbb{R}^{n}$, an initial approximation $x^{(0)}\in\mathbb{R}^{n}$, the maximal number of iterations $m$, a subspace embedding $S\in\mathbb{R}^{s\times n}$ ($s>m$), left and right preconditioners $M_{L}$ and $M_{R}$, and truncation parameter $t$, $\texttt{tol}_{\tau}$ and $\texttt{tol}_{B}$ for the restarting strategy.
 0: A computed solution $x\in\mathbb{R}^{n}$ approximating the solution of $Ax=b$.
-1:  $r^{(0)}\leftarrow b-Ax^{(0)}$ and $g\leftarrow SM_{L}^{-1}r^{(0)}$.
-2:  $\tilde{x}^{(0)}\leftarrow Sx^{(0)}$.
-3:  $B_{1}\leftarrow r^{(0)}/\|r^{(0)}\rVert$.
+[…]
+[…]
+[…]
 4:for $i=1:m$ do
@@
 9:  Solve the triangular system $T_{1:i}y^{(i)}=U_{1:i}^{\top}g$ to obtain $y^{(i)}$.
-10:     $\tilde{Z}_{i}=SZ_{i}$ and $\tilde{x}^{(i)}\leftarrow\tilde{x}^{(0)}+\tilde{Z}_{1:i}y^{(i)}$.
+10:  $\tilde{Z}_{i}=SZ_{i}$ and $\tilde{d}^{(i)}=C_{1:i}y^{(i)}$.
 11:  if the stopping criterion is satisfied then
-12:        return  $x=x^{(i)}\leftarrow x^{(0)}+Z_{1:i}y^{(i)}$.
+12:   return $x=x^{(i)}\leftarrow x^{(0)}+d^{(i)}$ with $d^{(i)}=Z_{1:i}y^{(i)}$.
 13:  end if
-14:     $\tilde{\tau}_{i}\leftarrow\|\tilde{Z}_{1:i}\rVert\|y^{(i)}\rVert/\|\tilde{x}^{(i)}\rVert$
+14:  $\tilde{\tau}_{i}\leftarrow\|\tilde{Z}_{1:i}\rVert\bigl\|A\bigr\rVert_{\mathsf{F}}\|y^{(i)}\rVert/\|\tilde{d}^{(i)}\rVert$.
 15:  if $\texttt{tol}_{\tau}\cdot\tilde{\tau}_{i}\geq 1$ then
-16:        if $\texttt{tol}_{B}\cdot\kappa(\tilde{Z}_{1:i})\geq 1$ then
-17:           $t\leftarrow 2t$
+16:   if $\tilde{\tau}_{i}>1.1\cdot\tilde{\tau}_{i-1}$ then
+17:    $t\leftarrow\min\{i+1,2t\}$.
 18:   end if
-19:        return
-20:     end if
-21:     if i == m then
-22:        $t\leftarrow 2t$
-23:     end if
-24:  end for
+19:  end if
+20:end for
 Algorithm 4 One cycle of preconditioned sketched GMRES (sGMRES) with adaptive restarting
@@
 All experiments were performed in MATLAB R2023a on a Windows 11 HP laptop with an 11th Gen Intel(R) Core processor with 2.80 GHz and 8 GB of RAM.
-We used (with modifications) the open source code provided by Güttel and Simunec as part of their recent work [7].
+We used (with modifications) the open source code provided by Güttel and Simunec as part of their recent work [guttel2024sketch].
 Our code is available at https://github.com/burkel8/stability_sGMRES.
-The sketching operator used in all experiments is the subsampled random Hadamard transform (see, e.g.,  [2]).
-We consider solving a linear system $Ax=b$ with sGMRES using both the truncated Arnoldi process and the sketch-and-select Arnoldi method (sGMRES-trunc and sGMRES-saa, respectively, in the plots).
-Although a few variants of the sketch-and-select Arnoldi algorithm have been proposed in [7], we use the sGMRES-saa-pinv method, as it was shown to produce the best conditioned basis in their experiments.
-We also ran the experiments with sGMRES-saa-greedy and obtained similar results, and thus do not include this method here.
-We note that the different variations of sketch-and-select Arnoldi differ in how the $t$ vectors with which to orthogonalize against are chosen, and we refer to [7] for more details on these methods.
+The sketching operator used in all experiments is the subsampled random Hadamard transform (see, e.g.,  [balabanov2025randomized]).
+We consider solving a linear system $Ax=b$ with sGMRES using both the truncated Arnoldi process and the sketch-and-select Arnoldi method (sGMRES-trunc and sGMRES-ssa, respectively, in the plots).
+Although a few variants of the sketch-and-select Arnoldi algorithm have been proposed in [guttel2024sketch], we use the sGMRES-ssa-pinv method, as it was shown to produce the best conditioned basis in their experiments.
+We also ran the experiments with sGMRES-ssa-greedy and obtained similar results, and thus do not include this method here.
+We note that the different variations of sketch-and-select Arnoldi differ in how the $t$ vectors with which to orthogonalize against are chosen, and we refer to [guttel2024sketch] for more details on these methods.
 In these experiments, we always choose the same value of $t$ in the sketch-and-select Arnoldi algorithm that we use as the truncation parameter in the truncated Arnoldi method.
@@
 In this subsection, we use three different examples to illustrate the results of our analysis shown in Sections 4 and 5.
-For all sketching techniques we take $s=2(m+1)$, where $m$ is the maximum length of each Arnoldi cycle, as suggested in [10].
+For all sketching techniques we take $s=2(m+1)$, where $m$ is the maximum length of each Arnoldi cycle, as suggested in [NT2024].
 We display the backward error, $\tau_{i}$ from (44), and the condition numbers of $B_{1:i}$ and $AB_{1:i}$ (approximated using the condition numbers of $SB_{1:i}$ and $SAB_{1:i}$, respectively) at each iteration.
@@
 The results are shown in Figure 1, where we see that the condition number of $B_{1:i}$ and $\tau_{i}$ generated by the truncated Arnoldi grows to $10^{16}$, leading to a poor backward error result for sGMRES-trunc.
-In the case of sGMRES-saa, the backward error is approximately $10^{-7}$ as both the condition number of $B_{1:i}$ and $\tau_{i}$ are not too large.
+In the case of sGMRES-ssa, the backward error is approximately $10^{-7}$ as both the condition number of $B_{1:i}$ and $\tau_{i}$ are not too large.
 For GMRES, it can be seen that the condition number of $B_{1:i}$ remains small, as does $\tau_{i}$, yielding a very small backward error after $400$ iterations.
@@
 We use this example to demonstrate that the bound outlined in Lemma 2 is sharper than the one mentioned in Theorem 1, and proves useful in certain situations.
-As shown in Figure 2, the behavior of the backward error for both sGMRES-trunc and sGMRES-saa diverges; notably, while the backward error for sGMRES-saa stops improving after about 160 iterations, sGMRES-trunc continues to reduce its error reaching a small level, despite both methods having ill-conditioned $B_{1:i}$.
+As shown in Figure 2, the behavior of the backward error for both sGMRES-trunc and sGMRES-ssa diverges; notably, while the backward error for sGMRES-ssa stops improving after about 160 iterations, sGMRES-trunc continues to reduce its error reaching a small level, despite both methods having ill-conditioned $B_{1:i}$.
 This indicates that for this scenario, the bounds shown by Theorem 1, specifically (29) and (51), are overly pessimistic and fail to account for the discrepancy observed.
-We further notice that $\tau_{i}$ for sGMRES-trunc is considerably smaller than for sGMRES-saa, explaining the superior backward error performance of sGMRES-trunc, as anticipated by Lemma 2.
+We further notice that $\tau_{i}$ for sGMRES-trunc is considerably smaller than for sGMRES-ssa, explaining the superior backward error performance of sGMRES-trunc, as anticipated by Lemma 2.
+We also observe but do not find a clear theoretical explanation for the phenomenon that $\tau_{i}$ can even decrease after $\kappa(AB)$ and $\kappa(B)$ are both very ill-conditioned, which means that solving the least squares problem is totally unstable.
 Figure 2: Example 2: From the left to right, the plots show the backward error, $\tau_{i}$ and the approximate condition numbers of the $B_{i}$ and $AB_{i}$, monitored every iteration, for different implementations of sGMRES, and an implementation of standard GMRES.
 6.1.3 Comparison between different sGMRES variants with and without restarting
-In the third example, the matrix $A$ is the Norris/stomach matrix of size $n=213,360$ obtained from the SuiteSparse matrix collection [4] (also tested in [7]).
+In the third example, the matrix $A$ is the Norris/stomach matrix of size $n=213,360$ obtained from the SuiteSparse matrix collection [Davis2011] (also tested in [guttel2024sketch]).
 The right-hand side vector is randomly generated according to the standard normal distribution.
@@
 Figure 3: Example 3: From the left to right, the plots show the backward error, $\tau_{i}$, and the approximate condition numbers of $B_{i}$ and $AB_{i}$, monitored every iteration, for different implementations of sGMRES, and an implementation of standard GMRES.
-6.2 sGMRES with the adaptively restarting strategy
-In the final experiment, we test the various adaptive restarting strategies discussed in Section 3. In particular, we test
-  • sGMRES (monitor $\tau_{i}$): monitor $\tau_{i}$, and when it gets too large, restart.
-  • sGMRES (monitor $\kappa(B_{1:i})$): monitor the condition number of $B_{1:i}$, and when it gets too large, restart.
-  • sGMRES (adapt monitor $\tau_{i}$): monitor $\tau_{i}$, and when it gets too large, double the truncation parameter $t$, and restart.
-Also restart and double $t$ every $m$ iterations even when $\tau_{i}$ is small.
-  • sGMRES (adapt monitor $\kappa(B_{1:i})$): monitor the condition number of $B_{1:i}$, and when it gets too large, double the truncation parameter $t$, and restart.
-Also restart and double $t$ every $m$ iterations even when the condition number of $B_{1:i}$ is small.
-  • sGMRES (adapt monitor $\tau_{i}$ and $\kappa(B_{1:i})$): monitor $\tau_{i}$, and when it gets too large, check the size of $\kappa(B_{1:i})$.
-If $\kappa(B_{1:i})$ is large, double the truncation parameter $t$, and restart.
-If $\kappa(B_{1:i})$ is not large enough, just restart without increasing $t$ (this method is shown in Algorithm 3).
-Also restart and double $t$ every $m$ iterations even when the condition number of $B_{1:i}$ is small.
+6.2 sGMRES with the adaptive restarting strategy
+In the final experiment, we test the adaptive restarting strategy discussed in Section 3. In particular, we test
   • restarted GMRES: restart every $m$ iterations.
-We solve a linear system by restarted sGMRES with the truncated Arnoldi process using all the various restarting strategies (beginning with $t=2$ for each method).
-We then plot the backward error curves produced by sGMRES with each restarting strategy, alongside the backward error curves produced by standard restarted GMRES.
-In each example, the right-hand side is randomly generated according to the standard normal distribution, and both $\texttt{tol}_{\tau}$ and $\texttt{tol}_{B}$ are taken to be the MATLAB machine precision obtained by the eps command.
-For GMRES, we restart a maximum of $d$ times, every $m$ iterations.
-We then allow a maximum of $md$ iterations for all implementations of sGMRES methods, and each implementation to restart according to its own strategy.
-In the first example, the matrix $A$ is the Norris/stomach matrix, $m=150$ and $d=4$.
-In the second example, the matrix $A$ is the fs_760_1 matrix obtained from SuiteSparse, and $m=30$ and $d=20$.
-In the final example, the matrix $A$ is taken to be the sherman2 matrix from SuiteSparse, with $m=200$ and $d=5$.
-The results are shown in Figure 4.
-The curve for sGMRES (monitor $\tau_{i}$) is hidden behind the curve for sGMRES (adapt monitor $\tau_{i}$) for each of the examples we tested.
-This is because $\tau_{i}$ is not directly dependent on the condition number of $B_{1:i}$, and so increasing $t$ does not help in this case.
-From the examples we have tested, sGMRES (adapt monitor $\tau_{i}$ and $\kappa(B_{1:i})$) seems to be the best approach as it produces backward error curves close to standard restarted GMRES in each of the examples.
+  • restarted sGMRES: restart every $m$ iterations with fixed parameter $t=1$.
+  • restarted sGMRES with adapt $t$: restart every $m$ iterations with adaptive parameter $t$ beginning with $t=1$ and $\texttt{tol}_{\tau}=\bm{u}$, i.e., Algorithm 4.
+Figures 4 and 5 illustrate tests of three algorithms on matrices with different condition numbers $5.49\times 10^{3}$ and $9.64\times 10^{11}$.In each example, the right-hand side is generated randomly following the standard normal distribution, with $m$ values of $50$ and $100$, and $\texttt{nrestarts}=10^{3}/m$.
+In Figure 4, both restarted GMRES and sGMRES achieve small backward errors, with restarted sGMRES using the adaptive restarting strategy accelerating convergence compared to a fixed $t$ approach by using a slightly larger $t$.
+As shown in the middle subfigure of Figure 4, it can be expected that the performance of restarted sGMRES, using an adaptive restarting strategy, tends towards the performance of restarted sGMRES.
+Figure 5 clearly shows that, for $m=100$, our adaptive restarting strategy can enhance the poor backward error that arises from a small $t$ by adaptively increasing $t$.
+This effect is even more pronounced for $m=400$.
+Conversely, when $m=50$, the backward error is limited by the small $m$, not $t$, since the behavior of restarted GMRES mirrors that of restarted sGMRES with a small $t$, which does not offer any improvement for this scenario.
+In this case, $t$ is repeatedly doubled up to $m$ during the early stage, thus restarted sGMRES with the adaptive strategy performs similarly to standard GMRES.
+By comparing the experiments shown in Figures 4 and 5, we observe that restarted GMRES also requires a large $m$ in order to obtain a small backward error.
+According to Theorem 2, the reduced factor $\Lambda^{(i)}_{1}$ defined in (46) depends on $\bar{\tau}_{k^{(i)}}$, whose upper bound is determined by $\kappa(A)$.
+Therefore, for well-conditioned matrices such as fs_760_1, achieving $\Lambda^{(i)}_{1}\ll 1$ is easy, meaning that a relatively small $m$ typically suffices.
+In contrast, for ill-conditioned matrices, a significantly larger $m$ may be necessary.
+Furthermore, in Figure 6, we combine the preconditioners generated by the MATLAB command ilu with the adaptive strategy to test sherman2, i.e., the same matrix tested in Figure 5.
+Due to the efficiency of the preconditioners, we can reduce the condition number of the preconditioned system and then employ much smaller $m$, that is, $m=5$, $10$.
+It can be seen that when preconditioning is used, restarted sGMRES with and without the adaptive strategy have similar behavior and require fewer orthogonalizations than standard GMRES.
+In general, sGMRES with our adaptive restarting approach often performs comparably to standard GMRES regarding the accuracy, and has the potential to result in substantial time savings especially for the well-conditioned case.
 Figure 4: Adaptive restarting experiment.
-From left to right, the backward error plots for Norris/Stomach, fs_760_1, and sherman2.
-The curve for sGMRES (monitor $\tau_{i}$), i.e., the dark blue line, is hidden behind the curve for sGMRES (adapt monitor $\tau_{i}$), i.e., the yellow line, for each of the examples we tested.
-The curve for sGMRES (adpt monitor $\tau_{i}$ and $\kappa(B_{1:i})$), i.e., the green line, is almost hidden behind the curve for GMRES, i.e., the light blue one.
+From left to right, we plot the backward error, the value of $\bar{t}$, and the value of $\tau$ for fs_760_1, where $\bar{t}=\max(t,i)$ represents the number of columns against which $B_{i}$ needs to perform orthogonalization in Line 2 of Algorithm 3.
+Figure 5: Adaptive restarting experiment.
+From left to right, we plot the backward error, the value of $\bar{t}$, and the value of $\tau$ for sherman2, where $\bar{t}=\max(t,i)$ represents the number of columns against which $B_{i}$ needs to perform orthogonalization in Line 2 of Algorithm 3.
+Figure 6: Adaptive restarting experiment with preconditioning.
+From left to right, we plot the backward error, the value of $\bar{t}$, and the value of $\tau$ for sherman2, where $\bar{t}=\max(t,i)$ represents the number of columns against which $B_{i}$ needs to perform orthogonalization in Line 2 of Algorithm 3.
 7 Conclusions
 In this paper, we performed a backward stability analysis of sketched GMRES and showed how the backward error is largely influenced by the condition number of the generated Krylov basis.
-However, in [7, Figure 9], it was demonstrated that the backward error of sGMRES can still decrease even when the basis becomes ill-conditioned.
-To provide more insight into this finding from [7], we proposed a sharper bound that relies on $\|\hat{B}_{1:k}\rVert\|\hat{y}^{(k)}\rVert/\|\hat{x}^{(k)}\rVert$ rather than $\kappa(\hat{B}_{1:k})$.
+However, in [guttel2024sketch, Figure 9], it was demonstrated that the backward error of sGMRES can still decrease even when the basis becomes ill-conditioned.
+To provide more insight into this finding from [guttel2024sketch], we proposed a sharper bound that relies on $\|\hat{B}_{1:k}\rVert\|\hat{y}^{(k)}\rVert/\|\hat{x}^{(k)}\rVert$ rather than $\kappa(\hat{B}_{1:k})$.
 We then showed how restarting can be used to stabilize a sketched GMRES implementation, and proposed an adaptive restarting strategy for sGMRES based on our analysis.
@@
 Neither the European Union nor the granting authority can be held responsible for them.
+L.
+Burke acknowledges funding from the InterFLOP (ANR-20-CE46-0009), NumPEx ExaMA (ANR-22-EXNU-0002), MixHPC (ANR-23-CE46-0005-01), and FPT-4 (ANR-24-CE46- 7572) projects of the French National Agency for Research (ANR).

### v3

@@
 Under this condition, we demonstrate that the stability of the sketched GMRES is mainly affected by the condition number of $B_{1:i}$, which can be mitigated by using the restarting technique.
-We also derive sharper bounds that explain why the backward error can be small even in cases when the basis $B_{1:i}$ is very ill-conditioned, which has been observed in the literature but not yet explained theoretically.
+We also derive sharper bounds that better capture the attainable backward error especially for cases when the basis $B_{1:i}$ is very ill-conditioned, which has been observed in the literature but not yet explained theoretically.
 We present numerical experiments to demonstrate the conclusions of our analysis, and also show that adaptively restarting where appropriate allows us to recover backward stability in sketched GMRES.
@@
 1 Introduction
-In this paper, we analyze the numerical stability of sketched GMRES (sGMRES) [NT2024] for solving nonsymmetric and nonsingular linear systems of the form
+In this paper, we analyze the numerical stability of sketched GMRES (sGMRES) [10] for solving nonsymmetric and nonsingular linear systems of the form
 $Ax=b,\kern 5.0ptA\in\mathbb{R}^{n\times n}$ (1)
 with a given matrix $A$ and a right-hand side $b$.
-Although standard GMRES [SaadSchultz1986] is one of the most popular methods to solve (1), it comes at the expense of constructing a matrix $V_{1:i}\in\mathbb{R}^{n\times i}$ with columns that form an orthonormal basis for the Krylov subspace
+Although standard GMRES [12] is one of the most popular methods to solve (1), it comes at the expense of constructing a matrix $V_{1:i}\in\mathbb{R}^{n\times i}$ with columns that form an orthonormal basis for the Krylov subspace
 $\mathcal{K}_{i}(A,r^{(0)})=\operatorname{span}\{r^{(0)},Ar^{(0)},\dotsc,A^{i-1}r^{(0)}\},$
 where $i$ is the iteration number, and $r^{(0)}=b-Ax^{(0)}$ is the initial residual corresponding to an initial approximation $x^{(0)}$.
-Sketched GMRES was proposed in [NT2024] as an attractive alternative to GMRES that exploits the technique of randomized sketching [martinsson2020randomized] to avoid the high cost of constructing an orthonormal basis.
+Sketched GMRES was proposed in [10] as an attractive alternative to GMRES that exploits the technique of randomized sketching [8] to avoid the high cost of constructing an orthonormal basis.
 Instead, the basis is only partially orthogonalized, thereby reducing the arithmetic cost from $O(ni^{2})$ to $O(i^{3}+ni\log i)$.
-It has been observed in numerical experiments in [NT2024] that sGMRES can fail once the condition number of the basis $B_{1:i}$ becomes too large, and methods such as the sketch-and-select Arnoldi algorithm [guttel2024sketch] have been proposed as a means to ensure the condition number of the basis $B_{1:i}$ does not grow too rapidly.
-Indeed, numerical experiments in [guttel2024sketch] showed that working with a well-conditioned basis improved the overall convergence behavior of sketched GMRES.
-In this paper, we rigorously analyze the backward stability of preconditioned sketched GMRES, and justify the experimental observations made in the literature [NT2024, guttel2024sketch].
+It has been observed in numerical experiments in [10] that sGMRES can fail once the condition number of the basis $B_{1:i}$ becomes too large, and methods such as the sketch-and-select Arnoldi algorithm [6] have been proposed as a means to ensure the condition number of the basis $B_{1:i}$ does not grow too rapidly.
+Indeed, numerical experiments in [6] showed that working with a well-conditioned basis improved the overall convergence behavior of sketched GMRES.
+In this paper, we rigorously analyze the backward stability of preconditioned sketched GMRES, and justify the experimental observations made in the literature [10, 6].
 Namely, we show that the backward error of sketched GMRES depends on $\kappa(Z_{1:i})$ as long as $\kappa(AZ_{1:i})$ is not too large, where $Z_{1:i}=M_{R}^{-1}B_{1:i}$ with the right preconditioner $M_{R}$.
@@
 The remainder of this paper is organized as follows.
-In Section 2, we give an overview of GMRES and the MOD-GMRES framework [BHMV2024] for performing a backward stability analysis of different GMRES variants.
+In Section 2, we give an overview of GMRES and the MOD-GMRES framework [3] for performing a backward stability analysis of different GMRES variants.
 We then discuss preconditioned sketched GMRES in Section 3, and analyze the backward stability of preconditioned sketched GMRES and restarted sketched GMRES, respectively, in Sections 4 and 5.
@@
 Furthermore, for optimal performance it is common to exploit techniques such as deflation, mixed precision, or randomization within GMRES.
-In order to understand the finite precision behavior of any given GMRES variant, a backward error analysis framework was developed in [BHMV2024], which simplifies the process of deriving bounds for the attainable normwise backward and forward errors of the computed solutions of different GMRES variants.
+In order to understand the finite precision behavior of any given GMRES variant, a backward error analysis framework was developed in [3], which simplifies the process of deriving bounds for the attainable normwise backward and forward errors of the computed solutions of different GMRES variants.
 The framework is built from a backward error analysis of a generic GMRES algorithm known as modular GMRES or MOD-GMRES, that is composed of three elemental operations:
@@
 By specializing these operations and meeting their assumptions, as outlined in the framework, modular GMRES can describe many GMRES implementations and variants, and can be used to perform their backward error analysis.
-For example, well-known results from [drkovsova1995numerical] were derived in [BHMV2024] using the framework.
+For example, well-known results from [5] were derived in [3] using the framework.
 In particular, it was shown that GMRES with Householder orthogonalization (HH-GMRES) produces a computed solution whose backward error is on the order of the unit roundoff, and is thus backward stable.
-Similarly, the backward stability result for GMRES with modified Gram-Schmidt (MGS-GMRES), derived in [paige2006modified], was also recovered using the framework in [BHMV2024].
+Similarly, the backward stability result for GMRES with modified Gram-Schmidt (MGS-GMRES), derived in [11], was also recovered using the framework in [3].
 3 Preconditioned Sketched GMRES
+Before introducing sketched GMRES, we first recall the basic definition of the random sketching technique.
+For $\epsilon\in(0,1)$, a sketching matrix $S\in\mathbb{R}^{s\times n}$ is an $\epsilon$-subspace embedding for the subspace $\mathcal{V}\subseteq\mathbb{R}^{n\times m}$ if $S$ satisfies that
+$(1-\epsilon)\|v\|^{2}\leq\|Sv\|^{2}\leq(1+\epsilon)\|v\|^{2},$ (4)
+for any $v\in\mathcal{V}$. In practice, the sketching matrix $S$ may not be explicitly available and needs to be drawn at random to achieve (4) with high probability.
 Sketched GMRES aims to avoid the $O(ni^{2})$ arithmetic cost of performing an expensive full orthogonalization of the basis vectors in standard GMRES.
@@
 The sketched problem (5) can then be constructed and solved inexpensively, even without orthonormality of the basis $B_{1:i}$.
-The solution of the sketched problem serves as a good approximation of the original problem as long as the matrix $S$ is chosen to be an $\epsilon$-subspace embedding for the Krylov subspace $\mathcal{K}_{i}(M_{L}^{-1}AM_{R}^{-1},M_{L}^{-1}r^{(0)})$, i.e., $S$ satisfies
-$(1-\epsilon)\|v\|^{2}\leq\|Sv\|^{2}\leq(1+\epsilon)\|v\|^{2},$ (5)
-for $\epsilon\in(0,1)$, and every $v\in\mathcal{K}_{i}(M_{L}^{-1}AM_{R}^{-1},M_{L}^{-1}r^{(0)})$.
-In practice, the sketching matrix $S$ may not be explicitly available and needs to be drawn at random to achieve (5) with high probability.
+The solution of the sketched problem serves as a good approximation of the original problem as long as the matrix $S$ is chosen to be an $\epsilon$-subspace embedding for the Krylov subspace $\mathcal{K}_{i}(M_{L}^{-1}AM_{R}^{-1},M_{L}^{-1}r^{(0)})$, i.e., $S$ satisfies (4) for $\epsilon\in(0,1)$, and every $v\in\mathcal{K}_{i}(M_{L}^{-1}AM_{R}^{-1},M_{L}^{-1}r^{(0)})$.
 Additionally, $S$ must be constructed without prior knowledge of the Krylov subspace $\mathcal{K}_{i}(M_{L}^{-1}AM_{R}^{-1},M_{L}^{-1}r^{(0)})$ and should typically be a valid embedding for a broad subspace, specifically $\mathbb{R}^{n\times i}$.
@@
 Algorithm 2 Restarted preconditioned sketched GMRES
-It is clear from experimental observations made in [NT2024, guttel2024sketch] that the basis condition number should not grow too large throughout the iteration in order to ensure numerical stability of sketched GMRES.
+It is clear from experimental observations made in [10, 6] that the basis condition number should not grow too large throughout the iteration in order to ensure numerical stability of sketched GMRES.
 The non-orthogonal basis can be constructed using the $t$ truncated Arnoldi process, as summarized in Algorithm 3, where each new Krylov vector is orthogonalized against the previous $t$ vectors.
-Alternatively, the sketch-and-select method [guttel2024sketch] uses the sketched basis $SB_{i}$ to select $t$ vectors, among all previously computed vectors, with which to perform the partial orthogonalization, with the aim of generating a basis with smaller condition number than that generated by truncated Arnoldi.
+Alternatively, the sketch-and-select method [6] uses the sketched basis $SB_{i}$ to select $t$ vectors, among all previously computed vectors, with which to perform the partial orthogonalization, with the aim of generating a basis with smaller condition number than that generated by truncated Arnoldi.
 0: A matrix $A\in\mathbb{R}^{n\times n}$, a truncated parameter $t$, the basis $B_{1:i-1}$ generated by the first i1i-1 truncated Arnoldi iterations.
@@
 4 Backward stability of preconditioned sketched GMRES
-In this section, we aim to analyze the backward stability of preconditioned sketched GMRES (Algorithm 1) by primarily following the analysis of the MOD-GMRES framework [BHMV2024].
+In this section, we aim to analyze the backward stability of preconditioned sketched GMRES (Algorithm 1) by primarily following the analysis of the MOD-GMRES framework [3].
 The first $i$ iterations of preconditioned sketched GMRES can be summarized in the following steps:
@@
 $\hat{W}_{1:i}=M_{L}^{-1}A\hat{Z}_{1:i}+\Delta W_{1:i},\quad\bigl\|\Delta W_{1:i}\bigr\rVert_{\mathsf{F}}\leq\varepsilon_{W}\|M_{L}^{-1}\rVert\bigl\|A\bigr\rVert_{\mathsf{F}}\bigl\|\hat{Z}_{1:i}\bigr\rVert_{\mathsf{F}},$ (7)
-with $\varepsilon_{W}\leq O(\bm{u})$ from the standard rounding error analysis of [H2002].
+with $\varepsilon_{W}\leq O(\bm{u})$ from the standard rounding error analysis of [7].
 Similarly, we assume that the computed result of $\tilde{b}:=M_{L}^{-1}b$ computed by Line 1 in Algorithm 1 satisfies
@@
 Step 2: Least squares solver
-In iteratively solving the least squares problem in Step 2, we assume that there is a special dimension $k\leq n$, known as the key dimension [BHMV2024], for which the computed solution is guaranteed to have reached a small backward error, i.e., $\hat{b}$ lies in the range of non-singular basis $\hat{W}_{1:k}$.
+In iteratively solving the least squares problem in Step 2, we assume that there is a special dimension $k\leq n$, known as the key dimension [3], for which the computed solution is guaranteed to have reached a small backward error, i.e., $\hat{b}$ lies in the range of non-singular basis $\hat{W}_{1:k}$.
 The definition of the key dimension is as follows:
@@
 $\begin{split}&\hat{y}^{(i)}=\operatorname*{\arg\min}_{y}\|\hat{g}+\Delta\hat{g}-(\hat{C}_{1:i}+\Delta\hat{C}_{1:i})y\rVert,\\ &\|\begin{bmatrix}\Delta\hat{g}&\Delta\hat{C}_{1:i}\end{bmatrix}e_{j}\rVert\leq\varepsilon_{sls}\|\begin{bmatrix}\hat{g}&\hat{C}_{1:i}\end{bmatrix}e_{j}\rVert,\quad\forall j\leq i+1,\end{split}$ (13)
-with $\varepsilon_{sls}\leq O(\bm{u})$ from [H2002, Theorem 20.3]. Substituting $\hat{C}_{1:i}$ and $\hat{g}$ in (13) with (11) and (12), we have
+with $\varepsilon_{sls}\leq O(\bm{u})$ from [7, Theorem 20.3]. Substituting $\hat{C}_{1:i}$ and $\hat{g}$ in (13) with (11) and (12), we have
 $\begin{split}&\hat{y}^{(i)}=\operatorname*{\arg\min}_{y}\big\|S\tilde{\hat{b}}+\Delta\tilde{g}-(S\hat{W}_{1:i}+\Delta\tilde{C}_{1:i})y\big\rVert,\\ &\|\Delta\tilde{g}\rVert\leq\|\Delta g\rVert+\|\Delta\hat{g}\rVert\leq(1+\varepsilon_{sls})\varepsilon_{S}\|S\hat{b}\rVert+\varepsilon_{sls}\|S\hat{b}\rVert\leq\bigl(\varepsilon_{sls}+(1+\varepsilon_{sls})\varepsilon_{S}\bigr)\|S\hat{b}\rVert,\\ &\|\Delta\tilde{C}_{j}\rVert\leq\|\Delta C_{j}\rVert+\|\Delta\hat{C}_{j}\rVert\leq\bigl(\varepsilon_{sls}+(1+\varepsilon_{sls})\varepsilon_{S}\bigr)\|S\hat{W}_{j}\rVert,\quad\forall j\leq i.\end{split}$ (14)
@@
 Proof.
-We follow the proof in [BHMV2024] and employ [paige2006modified, Theorem 2.4] to bound $\|r^{(k)}_{S}\rVert$. Applying [paige2006modified, Theorem 2.4] to the least squares problem (14), we have
+We follow the proof in [3] and employ [11, Theorem 2.4] to bound $\|r^{(k)}_{S}\rVert$. Applying [11, Theorem 2.4] to the least squares problem (14), we have
 $\|r^{(k)}_{S}\rVert^{2}\leq\sigma_{\min}^{2}\bigl(\begin{bmatrix}\phi(S\hat{b}+\Delta\tilde{g})&S\hat{W}_{1:k}+\Delta\tilde{C}_{1:k}\end{bmatrix}\bigr)\biggl(\frac{1}{\phi^{2}}+\frac{\|\hat{y}^{(k)}\rVert^{2}}{1-\delta_{k}^{2}}\biggr),$ (19)
@@
 $\delta_{k}=\frac{\sigma_{\min}\bigl(\begin{bmatrix}\phi(S\hat{b}+\Delta\tilde{g})&S\hat{W}_{1:k}+\Delta\tilde{C}_{1:k}\end{bmatrix}\bigr)}{\sigma_{\min}\bigl(S\hat{W}_{1:k}+\Delta\tilde{C}_{1:k}\bigr)}<1.$ (20)
-We choose $\phi$ such that $\phi^{-2}=\|\hat{y}^{(k)}\rVert^{2}/(1-\delta_{k}^{2})$ as in [BHMV2024, (3.17)], which can simplify (19) as
+We choose $\phi$ such that $\phi^{-2}=\|\hat{y}^{(k)}\rVert^{2}/(1-\delta_{k}^{2})$ as in [3, (3.17)], which can simplify (19) as
 $\|r^{(k)}_{S}\rVert\leq\sqrt{2}\,\sigma_{\min}\bigl(\begin{bmatrix}\phi(S\hat{b}+\Delta\tilde{g})&S\hat{W}_{1:k}+\Delta\tilde{C}_{1:k}\end{bmatrix}\bigr)\phi^{-1}.$ (21)
@@
 $\hat{x}^{(i)}=\hat{Z}_{1:i}\hat{y}^{(i)}+\Delta x^{(i)},\quad\|\Delta x^{(i)}\rVert\leq\varepsilon_{x}\|\hat{Z}_{1:i}\rVert\|\hat{y}^{(i)}\rVert$ (27)
-with $\varepsilon_{x}\leq O(\bm{u})$ from [H2002].
+with $\varepsilon_{x}\leq O(\bm{u})$ from [7].
 We are now ready to describe the backward stability of preconditioned sketched GMRES (Algorithm 1) in the following theorem.
 Theorem 1.
-Assume that Algorithm 1 is applied with a computed preconditioned basis $\hat{Z}_{1:i}\in\mathbb{R}^{n\times i}$ and an $\epsilon$-subspace embedding $S\in\mathbb{R}^{s\times n}$ satisfying (5) with high probability, where (7), (8), (11), (12), (13), and (27) are satisfied.
-Then there exists $k\leq n$ such that $c_{1}(n,k)\bigl(\varepsilon_{S}+\varepsilon_{sls}\bigr)\sqrt{\frac{1-\epsilon}{1+\epsilon}}\kappa(\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix})\geq 1$.
-If it further holds that
-$\sqrt{k}\,c_{2}(n,k)\bigl(\varepsilon_{S}+\varepsilon_{sls}\bigr)\sqrt{\frac{1+\epsilon}{1-\epsilon}}\kappa(\hat{W}_{1:k})\leq 1/2,\qquad\varepsilon_{x}\kappa(\hat{Z}_{1:k})\leq 1/2,$ (28)
+Assume that Algorithm 1 is applied with a computed preconditioned basis $\hat{Z}_{1:m}\in\mathbb{R}^{n\times m}$ and a randomized subspace embedding $S\in\mathbb{R}^{s\times n}$ satisfying (4) for the subspace $\mathbb{R}^{n\times m}$ with high probability, where (7), (8), (11), (12), (13), and (27) are satisfied.
+If there exists $k\leq m$ such that $c_{1}(n,k)\bigl(\varepsilon_{S}+\varepsilon_{sls}\bigr)\sqrt{\frac{1-\epsilon}{1+\epsilon}}\kappa(\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix})\geq 1$,
+$\sqrt{k}\,c_{2}(n,k)\bigl(\varepsilon_{S}+\varepsilon_{sls}\bigr)\sqrt{\frac{1+\epsilon}{1-\epsilon}}\kappa(\hat{W}_{1:k})\leq\frac{1}{2},\quad\text{and}\quad\varepsilon_{x}\kappa(\hat{Z}_{1:k})\leq\frac{1}{2},$ (28)
 then, with high probability,
@@
 Proof.
-Note that $\hat{b}\in\mathbb{R}^{n}$, and we can establish the existence of $k$, since the condition $c_{1}(n,k)\bigl(\varepsilon_{S}+\varepsilon_{sls}\bigr)\sqrt{\frac{1-\epsilon}{1+\epsilon}}\kappa(\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix})\geq 1$ is satisfied when $k=n$.
-Then we only need to prove (29). The residual of $\hat{x}^{(k)}$ can be bounded by the preconditioned residual as follows:
+The residual of $\hat{x}^{(k)}$ can be bounded by the preconditioned residual as follows:
 $\|b-A\hat{x}^{(k)}\rVert=\|M_{L}\bigl(M_{L}^{-1}b-M_{L}^{-1}A\hat{x}^{(k)})\bigr\rVert\leq\|M_{L}\rVert\|M_{L}^{-1}b-M_{L}^{-1}A\hat{x}^{(k)}\rVert,$ (31)
@@
 $\|M_{L}^{-1}b-M_{L}^{-1}A\hat{x}^{(k)}\rVert\leq\|\hat{b}-\hat{W}_{1:k}\hat{y}^{(k)}\rVert+\|\Delta W_{1:k}\hat{y}^{(k)}\rVert+\|\Delta b\rVert+\|M_{L}^{-1}\rVert\|A\Delta x^{(k)}\rVert.$ (32)
-Furthermore, from [NT2024, Fact 2.2] and the definition (15) of $r^{(k)}_{S}$, we have
+Furthermore, from [10, Fact 2.2] and the definition (15) of $r^{(k)}_{S}$, we have
 $\begin{split}\|\hat{b}-\hat{W}_{1:k}\hat{y}^{(k)}\rVert&\leq\frac{1}{1-\epsilon}\|S\hat{b}-S\hat{W}_{1:k}\hat{y}^{(k)}\rVert\\ &\leq\frac{1}{1-\epsilon}\|r^{(k)}_{S}\rVert+\frac{1}{1-\epsilon}\bigl(\|\Delta\tilde{g}\rVert+\bigl\|\Delta\tilde{C}_{1:k}\bigr\rVert_{\mathsf{F}}\|\hat{y}^{(k)}\rVert\bigr).\end{split}$ (33)
@@
 To apply Lemma 1 to bound $\|r^{(k)}_{S}\rVert$, we first need to verify (16) and (17).
-From [BG2022, Corollary 2.2], the condition numbers of the sketched matrices satisfy
+From [2, Corollary 2.2], the condition numbers of the sketched matrices satisfy
 $\kappa(S\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix})\geq\sqrt{\frac{1-\epsilon}{1+\epsilon}}\kappa(\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix})\quad\text{and}\quad\kappa(S\hat{W}_{1:k})\leq\sqrt{\frac{1+\epsilon}{1-\epsilon}}\kappa(\hat{W}_{1:k}),$
@@
 Theorem 1 indicates that (29) and (30) are satisfied with high probability.
-This is due to the $\epsilon$-subspace embedding, which generally meets the condition (5) with high probability.
+This is due to the choice of the subspace embedding $S$, which generally meets the condition (4) with high probability.
 Practically, the sketching method rarely fails, and slightly increasing the embedding dimension can further improve its reliability.
 In the worst case, $\epsilon$ may approach $1$ and the constant $(1+\epsilon)/(1-\epsilon)$ may be relatively large.
-For this reason, it is distracting to quantify the failure probabilities; see details in [MT2020].
+For this reason, it is distracting to quantify the failure probabilities; see details in [9].
+Remark 2.
+For standard GMRES, the existence of such a $k$ is straightforward to be guaranteed by observing that $\hat{b}\in\mathbb{R}^{n}$, because the condition $c_{1}(n,k)\bigl(\varepsilon_{S}+\varepsilon_{sls}\bigr)\sqrt{\frac{1-\epsilon}{1+\epsilon}}\,\kappa(\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix})\geq 1$ is automatically satisfied when $k=n$.
+In contrast, for sketched GMRES, one must select a proper subspace embedding $S\in\mathbb{R}^{s\times n}$ for the $m$-dimensional subspace in order to ensure the existence of such a $k$.
+When $s$ is sufficiently large—particularly in the extreme case s=ns=n—the existence of $k$ can be guaranteed in essentially the same way as for standard GMRES, but then the advantage of employing random sketching is lost.
 Remark 3.
@@
 from (7).
-Therefore, if $S$ is chosen to be a $n\times n$ identity matrix and $\hat{B}_{1:k}$ is chosen to be an orthonormal Krylov basis, Theorem 1 recovers a similar backward stability result as [BHMV2024, Theorem 3.1].
+Therefore, if $S$ is chosen to be a $n\times n$ identity matrix and $\hat{B}_{1:k}$ is chosen to be an orthonormal Krylov basis, Theorem 1 recovers a similar backward stability result as [3, Theorem 3.1].
 Notice that $\kappa(\hat{W}_{1:k})\leq 3\kappa(M_{L}^{-1}AZ_{1:j})$ from (41) if $O(\bm{u})\kappa(M_{L})\kappa(A)\kappa(\hat{Z}_{1:k})\leq 1$.
@@
 In particular, sGMRES is backward stable provided $\kappa(Z_{1:j})$ is not too large.
-It has been observed in numerical experiments in [guttel2024sketch, Figure 9] that in some cases, sGMRES will continue to converge even when $\kappa(\hat{Z}_{1:k})>10^{15}$, and so the bound shown in Theorem 1 is not tight.
+It has been observed in numerical experiments in [6, Figure 9] that in some cases, sGMRES will continue to converge even when $\kappa(\hat{Z}_{1:k})>10^{15}$, and so the bound shown in Theorem 1 is not tight.
 Therefore, we present the following lemma to show a tighter bound, which can be directly derived from (37) by noticing
@@
 Lemma 2.
-Assume that Algorithm 1 is applied with a computed preconditioned basis $\hat{Z}_{1:k}\in\mathbb{R}^{n\times k}$ and an $\epsilon$-subspace embedding $S\in\mathbb{R}^{s\times n}$, where (7), (8), (11), (12), (13), and (27) are satisfied.
-Then there exists $k\leq n$ such that $c_{1}(n,k)\bigl(\varepsilon_{S}+\varepsilon_{sls}\bigr)\sqrt{\frac{1-\epsilon}{1+\epsilon}}\kappa(\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix})\geq 1$.
-If it further holds that
-$\sqrt{k}\,c_{2}(n,k)\bigl(\varepsilon_{S}+\varepsilon_{sls}\bigr)\sqrt{\frac{1+\epsilon}{1-\epsilon}}\kappa(\hat{W}_{1:k})\leq 1/2,$
+Assume that Algorithm 1 is applied with a computed preconditioned basis $\hat{Z}_{1:m}\in\mathbb{R}^{n\times m}$ and a randomized subspace embedding $S\in\mathbb{R}^{s\times n}$ satisfying (4) for the subspace $\mathbb{R}^{n\times m}$ with high probability, where (7), (8), (11), (12), (13), and (27) are satisfied.
+If there exists $k\leq m$ such that $c_{1}(n,k)\bigl(\varepsilon_{S}+\varepsilon_{sls}\bigr)\sqrt{\frac{1-\epsilon}{1+\epsilon}}\kappa(\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix})\geq 1$,
+$\sqrt{k}\,c_{2}(n,k)\bigl(\varepsilon_{S}+\varepsilon_{sls}\bigr)\sqrt{\frac{1+\epsilon}{1-\epsilon}}\kappa(\hat{W}_{1:k})\leq\frac{1}{2},\quad\text{and}\quad\varepsilon_{x}\kappa(\hat{Z}_{1:k})\leq\frac{1}{2},$
 then, with high probability,
@@
 $\frac{\|b-A\hat{x}^{(k)}\rVert}{\bigl\|A\bigr\rVert_{\mathsf{F}}\|\hat{x}^{(k)}\rVert+\|b\rVert}\leq\frac{1+\epsilon}{1-\epsilon}O(\bm{u})\frac{\|\hat{B}_{1:k}\rVert\|\hat{y}^{(k)}\rVert}{\|\hat{x}^{(k)}\rVert}.$ (44)
-5 Adaptive restarting sGMRES
+5 Restarted sGMRES with adaptive truncation size
 Similar to the case of standard GMRES, restarting can be regarded as an iterative refinement process as follows:
@@
     Update the solution $x^{(i+1)}=x^{(i)}+d^{(i)}$.
-Furthermore, it can remove the influence of both $\kappa(M_{L})$ and $\kappa(Z_{1:j})$ on the backward stability of sGMRES, as shown in the following version of [BHMV2024, Theorem 4.1] describing the stability of restarted sketched GMRES (Algorithm 2).
+Furthermore, it can remove the influence of both $\kappa(M_{L})$ and $\kappa(Z_{1:j})$ on the backward stability of sGMRES, as shown in the following version of [3, Theorem 4.1] describing the stability of restarted sketched GMRES (Algorithm 2).
 Theorem 2.
-Assume that Algorithm 1 is restarted, such that for restart number $i\geq 1$, we compute a basis $\hat{Z}_{1:k^{(i)}}^{(i)}\in\mathbb{R}^{n\times k^{(i)}}$ and a randomized subspace embedding $S\in\mathbb{R}^{s\times n}$ satisfying (5) with high probability, where (7), (8), (11), (12), (13), and (27) are satisfied for $\varepsilon_{W}^{(i)}$, $\varepsilon_{sls}^{(i)}$, $\varepsilon_{S}^{(i)}$, $\varepsilon_{x}^{(i)}\leq O(\bm{u})$.
-Then there exists $k^{(i)}\leq n$ such that
-$O(\bm{u})\sqrt{\frac{1-\epsilon}{1+\epsilon}}\kappa\bigl(\begin{bmatrix}\phi r^{(i)}&\hat{W}_{1:k^{(i)}}^{(i)}\end{bmatrix}\bigr)\geq 1.$ (45)
-Assume that
-$\Lambda^{(i)}_{1}=\frac{1+\epsilon}{1-\epsilon}O(\bm{u})\kappa(M_{L})\bigl(1+\bar{\tau}_{k^{(i)}}\bigr)\ll 1$ (46)
+Assume that Algorithm 1 is restarted every $m$ iterations with a randomized subspace embedding $S\in\mathbb{R}^{s\times n}$ satisfying (4) for the subspace $\mathbb{R}^{n\times m}$ with high probability.
+Also, assume that for each restart index $i\geq 1$ there exists $k^{(i)}\leq m$ such that the computed preconditioned basis $\hat{Z}_{1:k^{(i)}}^{(i)}\in\mathbb{R}^{n\times k^{(i)}}$ satisfies (7), (8), (11), (12), (13), (27) with $\varepsilon_{W}^{(i)}$, $\varepsilon_{sls}^{(i)}$, $\varepsilon_{S}^{(i)}$, $\varepsilon_{x}^{(i)}\leq O(\bm{u})$, and that
+$\begin{split}&O(\bm{u})\sqrt{\frac{1-\epsilon}{1+\epsilon}}\kappa\bigl(\begin{bmatrix}\phi r^{(i)}&\hat{W}_{1:k^{(i)}}^{(i)}\end{bmatrix}\bigr)\geq 1,\qquad O(\bm{u})\sqrt{\frac{1+\epsilon}{1-\epsilon}}\kappa(\hat{W}_{1:k^{(i)}}^{(i)})\leq\frac{1}{2},\\ &\text{and}\quad O(\bm{u})\kappa(\hat{Z}_{1:k^{(i)}}^{(i)})\leq\frac{1}{2}.\end{split}$ (45)
+If
+$\Lambda^{(i)}_{1}=\frac{1+\epsilon}{1-\epsilon}O(\bm{u})\kappa(M_{L})\bigl(1+\bar{\tau}_{k^{(i)}}\bigr)\ll 1,$ (46)
 holds for $i$ with
-$\bar{\tau}_{k^{(i)}}=\frac{\|\hat{Z}^{(i)}_{1:k^{(i)}}\rVert\bigl\|A\bigr\rVert_{\mathsf{F}}\|\hat{y}^{(k^{(i)})}\rVert}{\|A\hat{d}^{(i)}\rVert}.$
+$\bar{\tau}_{k^{(i)}}=\frac{\|\hat{Z}^{(i)}_{1:k^{(i)}}\rVert\bigl\|A\bigr\rVert_{\mathsf{F}}\|\hat{y}^{(k^{(i)})}\rVert}{\|A\hat{d}^{(i)}\rVert},$
 then, with high probability, the backward error is reduced in the $i$-th restart cycle by a factor $\Lambda^{(i)}_{1}$ until it holds that
@@
 Proof.
-Similar to the proof of [BHMV2024, Theorem 4.1], we will employ [BHMV2024, Lemma 4.2], which requires verifying the assumption [BHMV2024, (4.14)], i.e., bounding $\|\hat{r}^{(i)}-A\hat{d}^{(i)}\rVert$ by $\|b-A\hat{x}^{(i)}\rVert$ and $\|b\rVert+\bigl\|A\bigr\rVert_{\mathsf{F}}\|\hat{x}^{(i)}\rVert$.
+Similar to the proof of [3, Theorem 4.1], we will employ [3, Lemma 4.2], which requires verifying the assumption [3, (4.14)], i.e., bounding $\|\hat{r}^{(i)}-A\hat{d}^{(i)}\rVert$ by $\|b-A\hat{x}^{(i)}\rVert$ and $\|b\rVert+\bigl\|A\bigr\rVert_{\mathsf{F}}\|\hat{x}^{(i)}\rVert$.
 From Lemma 2, we have
@@
 $\begin{split}\|\hat{r}^{(i)}-A\hat{d}^{(i)}\rVert&\leq\frac{1+\epsilon}{1-\epsilon}O(\bm{u})\kappa(M_{L})\bigl(1+\bar{\tau}_{k^{(i)}}\bigr)\|\hat{r}^{(i)}\rVert\\ &\leq\frac{1+\epsilon}{1-\epsilon}O(\bm{u})\kappa(M_{L})\bigl(1+\bar{\tau}_{k^{(i)}}\bigr)\bigl(\|b-A\hat{x}^{(i)}\rVert\\ &\quad+O(\bm{u})(\|b\rVert+\bigl\|A\bigr\rVert_{\mathsf{F}}\|\hat{x}^{(i)}\rVert)\bigr),\end{split}$ (49)
-which gives [BHMV2024, (4.14)]. Together with [BHMV2024, Lemma 4.2], we can draw the conclusion. ∎
+which gives [3, (4.14)]. Together with [3, Lemma 4.2], we can draw the conclusion. ∎
 Remark 7.
@@
 then, with high probability, the backward error is reduced in the $i$-th restart cycle by a factor $\Lambda^{(i)}_{1}$ until it holds that
-$\frac{\|b-A\hat{x}\rVert}{\bigl\|A\bigr\rVert_{\mathsf{F}}\|\hat{x}\rVert+\|b\rVert}\leq O(\bm{u}).$ (51)
+$\frac{\|b-A\hat{x}\rVert}{\bigl\|A\bigr\rVert_{\mathsf{F}}\|\hat{x}\rVert+\|b\rVert}\leq O(\bm{u}).$
 Remark 8.
@@
 Employing a small $t$ can result in considerable acceleration, though it might also introduce instability issues.
-Next, our aim is to propose an adaptive restarting strategy to recover from the instability of sGMRES caused by using a small $t$.
+Next, our aim is to propose an adaptive strategy to recover from the instability of sGMRES caused by using a small $t$.
 The definition of $\Lambda_{1}^{(1)}$ in Theorem 2 relying on $\bar{\tau}_{k^{(i)}}$, which influences the reduced factor of the backward error in the $i$-th restart cycle, implies that care should be taken to ensure $\bar{\tau}_{k^{(i)}}$ does not get too large as the iteration progresses.
-In fact, as noted in [NT2024], due to the subspace embedding property (5), $\bar{\tau}_{i}$ can be inexpensively estimated as
+In fact, as noted in [10], due to the subspace embedding property (4), $\bar{\tau}_{i}$ can be inexpensively estimated as
 $\tilde{\tau}_{i}:=\frac{\|SZ_{1:i}\rVert\bigl\|A\bigr\rVert_{\mathsf{F}}\|y^{(i)}\rVert}{\|SAZ_{1:i}y^{(i)}\rVert}.$
@@
 20:end for
-Algorithm 4 One cycle of preconditioned sketched GMRES (sGMRES) with adaptive restarting
+Algorithm 4 One cycle of preconditioned sketched GMRES (sGMRES) with adaptive truncation size
 6 Numerical Experiments
@@
 All experiments were performed in MATLAB R2023a on a Windows 11 HP laptop with an 11th Gen Intel(R) Core processor with 2.80 GHz and 8 GB of RAM.
-We used (with modifications) the open source code provided by Güttel and Simunec as part of their recent work [guttel2024sketch].
+We used (with modifications) the open source code provided by Güttel and Simunec as part of their recent work [6].
 Our code is available at https://github.com/burkel8/stability_sGMRES.
-The sketching operator used in all experiments is the subsampled random Hadamard transform (see, e.g.,  [balabanov2025randomized]).
+The sketching operator used in all experiments is the subsampled random Hadamard transform (see, e.g.,  [1]).
 We consider solving a linear system $Ax=b$ with sGMRES using both the truncated Arnoldi process and the sketch-and-select Arnoldi method (sGMRES-trunc and sGMRES-ssa, respectively, in the plots).
-Although a few variants of the sketch-and-select Arnoldi algorithm have been proposed in [guttel2024sketch], we use the sGMRES-ssa-pinv method, as it was shown to produce the best conditioned basis in their experiments.
+Although a few variants of the sketch-and-select Arnoldi algorithm have been proposed in [6], we use the sGMRES-ssa-pinv method, as it was shown to produce the best conditioned basis in their experiments.
 We also ran the experiments with sGMRES-ssa-greedy and obtained similar results, and thus do not include this method here.
-We note that the different variations of sketch-and-select Arnoldi differ in how the $t$ vectors with which to orthogonalize against are chosen, and we refer to [guttel2024sketch] for more details on these methods.
+We note that the different variations of sketch-and-select Arnoldi differ in how the $t$ vectors with which to orthogonalize against are chosen, and we refer to [6] for more details on these methods.
 In these experiments, we always choose the same value of $t$ in the sketch-and-select Arnoldi algorithm that we use as the truncation parameter in the truncated Arnoldi method.
-All comparisons to standard GMRES use GMRES with modified Gram-Schmidt.
-We do not use any preconditioner in the experiments, i.e., $M_{L}=M_{R}=I$.
-Finally, we drop the $\hat{.}$ notation in this section, as all quantities are computed quantities.
+All implementations of standard GMRES use the modified Gram-Schmidt Arnoldi process.
+We do not use any preconditioner, i.e., $M_{L}=M_{R}=I$, in the experiments except for Figure 6.
+Finally, we drop the $\hat{\cdot}$ notation in this section, as all quantities are computed quantities.
 6.1 Illustration of the backward error bound
 In this subsection, we use three different examples to illustrate the results of our analysis shown in Sections 4 and 5.
-For all sketching techniques we take $s=2(m+1)$, where $m$ is the maximum length of each Arnoldi cycle, as suggested in [NT2024].
+For all sketching techniques we take $s=2(m+1)$, where $m$ is the maximum length of each Arnoldi cycle, as suggested in [10].
 We display the backward error, $\tau_{i}$ from (44), and the condition numbers of $B_{1:i}$ and $AB_{1:i}$ (approximated using the condition numbers of $SB_{1:i}$ and $SAB_{1:i}$, respectively) at each iteration.
@@
 6.1.3 Comparison between different sGMRES variants with and without restarting
-In the third example, the matrix $A$ is the Norris/stomach matrix of size $n=213,360$ obtained from the SuiteSparse matrix collection [Davis2011] (also tested in [guttel2024sketch]).
+In the third example, the matrix $A$ is the Norris/stomach matrix of size $n=213,360$ obtained from the SuiteSparse matrix collection [4] (also tested in [6]).
 The right-hand side vector is randomly generated according to the standard normal distribution.
@@
 Figure 3: Example 3: From the left to right, the plots show the backward error, $\tau_{i}$, and the approximate condition numbers of $B_{i}$ and $AB_{i}$, monitored every iteration, for different implementations of sGMRES, and an implementation of standard GMRES.
-6.2 sGMRES with the adaptive restarting strategy
-In the final experiment, we test the adaptive restarting strategy discussed in Section 3. In particular, we test
+6.2 sGMRES with adaptive truncation size
+In the final experiment, we test the adaptive strategy discussed in Section 3. In particular, we test
   • restarted GMRES: restart every $m$ iterations.
@@
   • restarted sGMRES with adapt $t$: restart every $m$ iterations with adaptive parameter $t$ beginning with $t=1$ and $\texttt{tol}_{\tau}=\bm{u}$, i.e., Algorithm 4.
-Figures 4 and 5 illustrate tests of three algorithms on matrices with different condition numbers $5.49\times 10^{3}$ and $9.64\times 10^{11}$.In each example, the right-hand side is generated randomly following the standard normal distribution, with $m$ values of $50$ and $100$, and $\texttt{nrestarts}=10^{3}/m$.
-In Figure 4, both restarted GMRES and sGMRES achieve small backward errors, with restarted sGMRES using the adaptive restarting strategy accelerating convergence compared to a fixed $t$ approach by using a slightly larger $t$.
-As shown in the middle subfigure of Figure 4, it can be expected that the performance of restarted sGMRES, using an adaptive restarting strategy, tends towards the performance of restarted sGMRES.
-Figure 5 clearly shows that, for $m=100$, our adaptive restarting strategy can enhance the poor backward error that arises from a small $t$ by adaptively increasing $t$.
+  • restarted GMRES with $m=t_{\max}$: restart every $t_{\max}$ iterations, where $t_{\max}$ denotes the maximum truncation size employed in the “restarted sGMRES with adapt tt” method.
+Figures 4 and 5 illustrate tests of these algorithms on matrices fs_760_1 and sherman2, whose condition numbers are $5.49\times 10^{3}$ and $9.64\times 10^{11}$, respectively.
+In each example, the right-hand side is generated randomly following the standard normal distribution, with $m$ values of $50$ and $100$, and $\texttt{nrestarts}=1.5\cdot 10^{3}/m$.
+It should be noted that restarted GMRES with $m=t_{\max}$ is examined primarily to enable a direct comparison between restarted sGMRES with adaptive $t$ and restarted GMRES with similarly low orthogonalization costs.
+In Figure 5, we omit the result for “restarted GMRES with $m=t_{\max}$” because it coincides with that of “restarted GMRES”.
+In Figure 4, both restarted GMRES and sGMRES achieve small backward errors, with restarted sGMRES using the adaptive strategy accelerating convergence compared to a fixed $t$ approach by using a slightly larger $t$.
+As shown in the middle subfigure of Figure 4, it can be expected that the performance of restarted sGMRES, using adaptive truncation size, tends towards the performance of restarted sGMRES.
+To further illustrate the potential performance advantage of our adaptive strategy, we also compare the number of floating-point operations required for orthogonalization in Table 1.
+From the table, we observe that restarted sGMRES, both with and without the adaptive strategy, entails a comparable but substantially lower orthogonalization cost than restarted GMRES.
+Figure 4: Adaptive truncation size experiment.
+From left to right, we plot the backward error, the value of $\bar{t}$, and the value of $\tau$ for fs_760_1, where $\bar{t}=\max(t,i)$ represents the number of columns against which $B_{i}$ needs to perform orthogonalization in Line 2 of Algorithm 3 and $t_{\max}$ denotes the maximum truncation size employed in the “restarted sGMRES with adapt tt” method.
+We next present a difficult example for restarted sGMRES in Figure 5.
+In this situation, restarted sGMRES struggles to attain a small backward error.
+As indicated in Theorem 2, the attainable backward error is affected by two conditions, namely (45) and (46).
+For $m=50$ in Figure 5, the value of $m$ is too small for (45) to be satisfied.
+Consequently, even if we use restarted sGMRES with the adaptive strategy or restarted GMRES, both of which ensure a small $\tau$ so that (46) holds, restarting still fails to improve the attainable backward error.
+In contrast, when $m=100$ and $m=400$, the larger values of $m$ allow (45) to be satisfied.
+In this situation, $\tau$, associated with (46), becomes the dominant factor determining the attainable backward error.
+Figure 5 clearly shows that, for $m=100$, our adaptive strategy can enhance the poor backward error that arises from a small $t$ by adaptively increasing $t$ to reduce $\tau$ such that (46) is satisfied.
 This effect is even more pronounced for $m=400$.
-Conversely, when $m=50$, the backward error is limited by the small $m$, not $t$, since the behavior of restarted GMRES mirrors that of restarted sGMRES with a small $t$, which does not offer any improvement for this scenario.
-In this case, $t$ is repeatedly doubled up to $m$ during the early stage, thus restarted sGMRES with the adaptive strategy performs similarly to standard GMRES.
+In this case, $t$ is repeatedly doubled up to $m$ during the early stage, thus restarted sGMRES with the adaptive strategy utilizes a comparable orthogonalization cost as standard GMRES.
 By comparing the experiments shown in Figures 4 and 5, we observe that restarted GMRES also requires a large $m$ in order to obtain a small backward error.
-According to Theorem 2, the reduced factor $\Lambda^{(i)}_{1}$ defined in (46) depends on $\bar{\tau}_{k^{(i)}}$, whose upper bound is determined by $\kappa(A)$.
-Therefore, for well-conditioned matrices such as fs_760_1, achieving $\Lambda^{(i)}_{1}\ll 1$ is easy, meaning that a relatively small $m$ typically suffices.
-In contrast, for ill-conditioned matrices, a significantly larger $m$ may be necessary.
+According to Theorem 2, $O(\bm{u})\sqrt{\frac{1+\epsilon}{1-\epsilon}}\kappa(\hat{W}_{1:k^{(i)}}^{(i)})\leq 1/2$ in (45) depends on $\kappa(\hat{W}_{1:k^{(i)}}^{(i)})$, whose upper bound is determined by $\kappa(A)$ and $\kappa(\hat{Z}_{1:k^{(i)}}^{(i)})$.
+This implies that, in the worst case, it may be necessary to require $O(\bm{u})\kappa(\hat{Z}_{1:k^{(i)}}^{(i)})\leq 1/(2\kappa(A))$ in order to guarantee (45).
+Consequently, for well-conditioned matrices such as fs_760_1, $\kappa(\hat{Z}_{1:k^{(i)}}^{(i)})$ can be relatively large while still satisfying $O(\bm{u})\kappa(\hat{Z}_{1:k^{(i)}}^{(i)})\leq 1/(2\kappa(A))$, meaning that a relatively small truncation size usually suffices.
+In contrast, for ill-conditioned matrices, a significantly larger small truncation size may be required.
+Figure 5: Adaptive truncation size experiment.
+From left to right, we plot the backward error, the value of $\bar{t}$, and the value of $\tau$ for sherman2, where $\bar{t}=\max(t,i)$ represents the number of columns against which $B_{i}$ needs to perform orthogonalization in Line 2 of Algorithm 3.
 Furthermore, in Figure 6, we combine the preconditioners generated by the MATLAB command ilu with the adaptive strategy to test sherman2, i.e., the same matrix tested in Figure 5.
-Due to the efficiency of the preconditioners, we can reduce the condition number of the preconditioned system and then employ much smaller $m$, that is, $m=5$, $10$.
-It can be seen that when preconditioning is used, restarted sGMRES with and without the adaptive strategy have similar behavior and require fewer orthogonalizations than standard GMRES.
-In general, sGMRES with our adaptive restarting approach often performs comparably to standard GMRES regarding the accuracy, and has the potential to result in substantial time savings especially for the well-conditioned case.
-Figure 4: Adaptive restarting experiment.
-From left to right, we plot the backward error, the value of $\bar{t}$, and the value of $\tau$ for fs_760_1, where $\bar{t}=\max(t,i)$ represents the number of columns against which $B_{i}$ needs to perform orthogonalization in Line 2 of Algorithm 3.
-Figure 5: Adaptive restarting experiment.
-From left to right, we plot the backward error, the value of $\bar{t}$, and the value of $\tau$ for sherman2, where $\bar{t}=\max(t,i)$ represents the number of columns against which $B_{i}$ needs to perform orthogonalization in Line 2 of Algorithm 3.
-Figure 6: Adaptive restarting experiment with preconditioning.
-From left to right, we plot the backward error, the value of $\bar{t}$, and the value of $\tau$ for sherman2, where $\bar{t}=\max(t,i)$ represents the number of columns against which $B_{i}$ needs to perform orthogonalization in Line 2 of Algorithm 3.
+Due to the efficiency of the preconditioners, a small $m$, that is, $m=50$, is sufficient to achieve convergence.
+From Figure 6 together with Table 1, it can be seen that when preconditioning is used, restarted sGMRES with and without the adaptive strategy have similar behavior and require fewer orthogonalizations than standard GMRES.
+In general, sGMRES with adaptive truncation size often performs comparably to standard GMRES regarding the accuracy, and has the potential to result in substantial time savings especially for the well-conditioned case.
+Figure 6: Adaptive truncation size experiment with preconditioning.
+From left to right, we plot the backward error, the value of $\bar{t}$, and the value of $\tau$ for sherman2, where $\bar{t}=\max(t,i)$ represents the number of columns against which $B_{i}$ needs to perform orthogonalization in Line 2 of Algorithm 3 and $t_{\max}$ denotes the maximum truncation size employed in the “restarted sGMRES with adapt tt” method.
+Table 1: Comparison of the relative number of floating-point operations used in orthogonalization.
+In the table, we report the ratios of the number of floating-point operations required for the orthogonalization of “restarted sGMRES,” “restarted sGMRES with adapt tt,” and “restarted GMRES with $m=t_{\max}$” relative to that of “restarted GMRES,” respectively.
+m
+restarted
+GMRES
+restarted
+sGMRES
+restarted
+sGMRES
+with adapt $t$
+restarted
+GMRES
+with $m=t_{\max}$
+fs_760_1 without pre 50 1 0.3072 0.5959 1.7979
+100 1 0.5468 0.4485 3.4323
+fs_760_1
+with pre
+50 1 0.1250 0.1250 0.1883
 7 Conclusions
 In this paper, we performed a backward stability analysis of sketched GMRES and showed how the backward error is largely influenced by the condition number of the generated Krylov basis.
-However, in [guttel2024sketch, Figure 9], it was demonstrated that the backward error of sGMRES can still decrease even when the basis becomes ill-conditioned.
-To provide more insight into this finding from [guttel2024sketch], we proposed a sharper bound that relies on $\|\hat{B}_{1:k}\rVert\|\hat{y}^{(k)}\rVert/\|\hat{x}^{(k)}\rVert$ rather than $\kappa(\hat{B}_{1:k})$.
-We then showed how restarting can be used to stabilize a sketched GMRES implementation, and proposed an adaptive restarting strategy for sGMRES based on our analysis.
-Our numerical experiments indicate that the convergence of sketched GMRES with our adaptive restarting approach often performs comparably to standard GMRES and has the potential to result in substantial time savings.
+However, in [6, Figure 9], it was demonstrated that the backward error of sGMRES can still decrease even when the basis becomes ill-conditioned.
+To provide more insight into this finding from [6], we proposed a sharper bound that relies on $\|\hat{B}_{1:k}\rVert\|\hat{y}^{(k)}\rVert/\|\hat{x}^{(k)}\rVert$ rather than $\kappa(\hat{B}_{1:k})$.
+Nevertheless, it remains unclear why $\|\hat{B}_{1:k}\rVert\|\hat{y}^{(k)}\rVert/\|\hat{x}^{(k)}\rVert$ itself can even decrease—leading to a reduction in the backward error of sGMRES—when the basis becomes ill-conditioned.
+We then showed how restarting can be used to stabilize a sketched GMRES implementation, and proposed an adaptive strategy to choose proper truncation sizes for sGMRES based on our analysis.
+Our numerical experiments indicate that the convergence of sketched GMRES with adaptive truncation size often performs comparably to standard GMRES and has the potential to result in substantial time savings.
 Acknowledgments